### v1

# Deep Time Series Models: A Comprehensive Survey and Benchmark

Abstract
Time series, characterized by a sequence of data points arranged in a discrete-time order, are ubiquitous in real-world applications. Different from other modalities, time series present unique challenges due to their complex and dynamic nature, including the entanglement of nonlinear patterns and time-variant trends. Analyzing time series data is of great significance in real-world scenarios and has been widely studied over centuries. Recent years have witnessed remarkable breakthroughs in the time series community, with techniques shifting from traditional statistical methods to advanced deep learning models. In this paper, we delve into the design of deep time series models across various analysis tasks and review the existing literature from two perspectives: basic modules and model architectures. Further, we develop and release Time Series Library (TSLib) as a fair benchmark of deep time series models for diverse analysis tasks, which implements 24 mainstream models, covers 30 datasets from different domains, and supports five prevalent analysis tasks. Based on TSLib, we thoroughly evaluate 12 advanced deep time series models on different tasks. Empirical results indicate that models with specific structures are well-suited for distinct analytical tasks, which offers insights for research and adoption of deep time series models. Code is available at https://github.com/thuml/Time-Series-Library.
Index Terms:
Time series analysis, deep time series models, survey, benchmark1 Introduction
Time series refers to a sequence of data points indexed in a discrete-time order [1, 2], which are omnipresent in real-world applications, such as financial risk assessment, energy sustainability, and weather forecasting. Driven by the increasing availability of vast amounts of time series data across various domains, the community of time series analysis has witnessed tremendous advancements. Compared to image and text data, which have objectively prescribed syntax or intuitive patterns, the semantic information of time series data is primarily derived from the temporal variation [3]. This presents significant challenges in understanding the data, such as identifying sequential dependencies, trends, seasonal patterns, and complicated dynamics. Consequently, analyzing time series data requires sophisticated methods to capture and utilize these complex temporal representations.
Given the crucial role of time series data in real-world applications [4, 5, 6], time series analysis has been a longstanding research direction. Time series analysis encompasses the process of analyzing the temporal variation to understand time series data and make accurate predictions and informed decisions. One of the essential cornerstone of time series analysis is discovering the underlying patterns in time series data, which involves the intricate temporal dependencies and variate correlations inherent within the data. By capturing these complex dependencies, time series models can effectively reveal the underlying dynamics, and facilitate various downstream tasks, including forecasting, classification, imputation, and anomaly detection.
Traditional time series methods, such as AutoRegressive Integrated Moving Average (ARIMA) [1], Exponential Smoothing, and Spectral Analysis [7], have long served as stalwart tools in time series analysis. These models, grounded in statistical methodologies, have been instrumental in discovering patterns, trends, and seasonality within temporal variations. However, their capabilities are hindered due to the inherent limitations of capturing complex nonlinear relationships and long-term dependencies present in real-world time series data. The rigid assumptions of linearity and stationarity that underpin traditional models constrain their adaptability to eventful and evolving data flows.
| Survey | Analysis Task | Model Architecture | Benchmark | |||||||
| Forecasting | Classification | Imputation | Anomaly | MLP | CNN | RNN | GNN | Transformer | ||
| Fawaz et al. (2019) [8] | ✓ | ✓ | ✓ | ✓ | ||||||
| Braei et al. (2020) [9] | ✓ | ✓ | ✓ | ✓ | ||||||
| Torres et al. (2021) [10] | ✓ | ✓ | ✓ | ✓ | ||||||
| García et al. (2021) [11] | ✓ | ✓ | ✓ | ✓ | ||||||
| Wen et al. (2022) [12] | ✓ | ✓ | ✓ | ✓ | ||||||
| Jin et al. (2023) [13] | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
| Shao et al. (2023) [14] | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| Qiu et al. (2024) [15] | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| Our Survey | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
Deep models have garnered significant attention and achieved remarkable performance across various domains, including natural language processing (NLP) [16], [17], computer vision (CV) [18], [19], and recommendation systems [20]. In recent years, deep learning models [21, 22, 23, 24, 3] have demonstrated their capability to capture the intricate dependencies within time series data, making deep learning models a powerful tool for time series analysis over traditional statistical methods. More recently, Transformer models with attention mechanisms, originally developed for natural language processing tasks, have presented stunning power in processing large-scale data [17] and have also been adapted for learning time series data. These architectures offer the advantage of selectively focusing on different parts of the input sequence, allowing for more nuanced discovery of both temporal and variable dependencies in time series.
Related Surveys Although various time series models designed for different analysis tasks have emerged in recent years, there is a lack of a comprehensive overview of existing methods, covering both tasks and models. Previous reviews focus exclusively on either a specific model architecture or an analysis task. For example, [8, 9, 10, 11] reviews deep learning methods for specific time series analysis tasks while failing to include advanced architecture such as Transformer. Several surveys [12, 13] provide up-to-date reviews for time series analysis focusing on specific deep learning architectures(i.e., Graph Neural Network and Transformer). Recently, BasicTS [14] and TFB [15] introduce forecasting benchmarks that enable an unbiased evaluation of existing approaches but do not provide an overview of the architectural design of those deep models.
In this survey, we provide a comprehensive review of deep time series models for researchers and practitioners, starting from the basic modules to modern architectures. To foster practical applications, a time series benchmark is offered for a fair evaluation and identifying the effective scope of existing models. Our survey is organized as follows. Section 2 provides the background concepts of time series analysis. Section 3 introduces the basic modules that are widely utilized in prevalent deep time series models. Section 4 reviews the existing deep time series models in terms of the architecture design. Section 5 introduces the proposed open-source benchmark—Time Series Library (TSLib)—and presents extensive experimental comparison with detailed analysis. Section 6 provides a brief discussion of future research directions while Section 7 summarizes this survey.
2 Preliminaries
2.1 Time Series
Time series is a sequence of observations ordered by time, which can be denoted as , where represents the observed values at time point and is the number of variables. Since time series data are physical measurements obtained from sensors, systems are often recorded with multiple variables. Consequently, real-world time series usually recorded in a multivariate form. Theoretical studies [25, 26] have shown that when there are two or more non-stationary series, a linear combination of them can be stationary. This co-integration property helps in uncovering and modeling long-term relationships among non-stationary series. Therefore, the essence of time series analysis is to capture and utilize the temporal dependencies and inter-variable correlations within the observations.
Temporal Dependency Given the sequential nature inherent in the observations, one evident technological paradigm is to capture the temporal dependence of a set of historical data. The basic idea of temporal dependencies is the intricate correlations between time points or sub-series. Traditional statistical models have laid the groundwork for modeling temporal dependencies. Prominent models include ARIMA (Autoregressive Integrated Moving Average) [1] have been extensively studied for capturing complex temporal patterns in the time series modality. Owing to their simplicity and interpretability, these statistical methods remain popular for tasks where the underlying temporal dynamics do not exhibit high complexity. Considering the high-dimensionality and non-stationarity of real-world time series, the research focus shifted towards deep learning for time series analysis. These advanced methods are designed to handle more complex temporal dynamics and offer greater flexibility in capturing the temporal dependency of time series data.
Variate Correlation In addition to capturing temporal dependencies, understanding the variate correlations within high-dimensionality plays a pivotal role in analyzings multivariate time series. These correlations refer to the complex interactions and associations among different variables changing across the time. They provide valuable insights into the underlying dynamics and dependencies among the measurements, enabling a more comprehensive understanding of the latent process. Traditional approaches, such as Vector Autoregressive (VAR) models [27], extend the concept of autoregression to multiple variables and can capture the relationships between multiple quantities as they evolve over time. Technically, VAR represents each variable as a linear combination of its lagged values and the lagged values of all other variables in the model, which results in an inability to capture complex and non-linear relationships. Recently, advanced deep models, such as Graph Neural Networks [28] and Transformers [29, 30], have also been introduced for variate correlation modeling.
2.2 Time Series Analysis Tasks
Based on the understanding of underlying patterns and trends within time series data, time series analysis encompasses various downstream applications, including forecasting [31, 32], imputation [33, 34, 35], classification [8, 36], and anomaly detection [37, 5], each serving distinct purposes in diverse application domains.
We illustrate representative time series analysis tasks in Figure 1. Forecasting is a fundamental task in time series analysis that requires models to uncover temporal dependencies and dynamic patterns in the data. By capturing the relationships between past and future data, the forecasting model aims to predict future values or trends of the input series. Missing data due to sensor failures, data corruption, or absent measurements are ubiquitous in practical applications, leading to a growing demand for time series imputation to obtain higher-quality data. Unlike forecasting, which predicts future values based on historical observations, imputation focuses on reconstructing missing values using the available contextual information. Anomaly detection involves identifying unusual or abnormal patterns within a time series, which can indicate critical events, system faults, or outliers requiring further investigation. Lastly, classification assigns a label or category to a given time series based on its characteristics, a task widely utilized in fields such as medical diagnosis.
3 Basic Modules
Time series modeling approaches have evolved significantly, transitioning from traditional statistical models to sophisticated deep learning models. Despite these advancements, many classical tools and analytical algorithms remain widely used and continue to serve as foundational design principles in modern deep models. In this section, we focus on the major tools of classical time series analysis and demonstrate how they have been integrated as fundamental components in contemporary deep time series models.
3.1 Stationarization
As a foundational concept in time series analysis, stationarity refers to the property of a time series where its statistical properties remain constant over time. A stationary time series has a constant mean and variance, which simplifies statistical analysis and makes it easier to capture the underlying patterns and behavior within a time series. Since many statistics-based time series analysis methods take stationarity as a basic assumption, stationarization of time series data has become an essential module. There are ways of transforming non-stationary time series into stationary. Traditional time series models stationarize the time series through differencing or log-transformation. In recent deep learning approaches, data normalization [38] takes the role of stationarization in a simple but effective way, which standardizes the value distribution of observations while maintaining the intrinsic variations and further helps mitigate the distribution shift between the source and target domains.
The deep adaptive input normalization (DAIN) layer [39] was proposed to adaptively stationarize time series data according to their original distribution. RevIN [40] introduces reversible instance normalization to time series data, which is an effective normalization-and-denormalization method with learnable affine transforms to make the model bypass the non-stationary inputs. Non-Stationary Transformer [41] (Stationary for short in the following) proposes a simpler but more effective series stationarization technique that improves the predictive capability of non-stationary series without extra parameters. Specifically, for a sequence with time stamps and variates , the outline of Stationary [41] can be summarized as:
where is in a small value for numerical stability. are the variate-specific mean and variance. To recover the distribution and non-stationarity of the original series, a de-normalization module is further used to augment the model output with mean and variance statistics of inputs.
The idea of stationarization and the above-mentioned techniques have been widely used in subsequent deep time series models [23, 30, 42, 43]. Recent SAN [44] rethinks the nature of non-stationary data and tries to split it into non-overlap equally-sized slices and perform normalization on each slice. Specifically, based on the evolving trends of statistical properties, SAN introduces a statistics prediction module to predict the distributions of future slices.
3.2 Decomposition
Decomposition [45, 46], as a conventional approach in time series analysis, can disentangle time series into several components with categorized patterns, and works primarily useful for exploring complex series variations. In the previous work, diverse decomposition paradigms are explored.
3.2.1 Seasonal-Trend Decomposition
Seasonal-trend decomposition [47, 48] is one of the most common practices to make raw data more predictable, which can separate the series into several different components: trend, seasonal, cyclical, and irregular, namely
where the trend component represents the overall long-term pattern of the data over time, the cyclical component reflects repeated but non-periodic fluctuations within data, the seasonal component indicates the repetitive patterns over a fixed period, and the irregular component is the residuals or remainder of the time series after the other components have been removed.
The trend-seasonality decomposition can be achieved by using mathematical tools such as filters or exponential smoothing [49, 50]. Previous statistical approaches mainly adopt the trend-seasonality decomposition as data pre-processing [51]. In deep models, Autoformer [22] firstly introduces the idea of decomposition to deep learning architecture and proposes a series decomposition block as a basic module to extract the seasonal and trend-cyclical parts of deep features and input series, whose computation process can be formalized as:
The series decomposition block is concisely implemented based on a temporally average pooling layer with padding operation to keep the sequence length unchanged. This design can capture trends , and the remainder is taken as the seasonal part . The proposed series decomposition block has been widely used in the follow-up [6, 52, 53, 54, 55] as a native building block of deep models to disentangle the underlying patterns of deep features.
3.2.2 Basis Expansion
Basis expansion is a mathematical method used to represent a function or a set of data points in terms of a new set of pre-defined functions. These new functions form a basis for a function space, meaning any function in that space can be expressed as a linear combination of these basis functions. In the context of time series analysis, basis expansion is used to reveal complex non-linear temporal relationships by decomposing the time series into a combination of basic variations, which also enhances interpretability. As a representative model, N-BEATS [56] presents hierarchical decomposition to time series by utilizing a fully connected layer to produce expansion coefficients for both backward and forward forecasts. For -th blocks in the proposed hierarchical architecture, the operation can be as follows:
where is the backcast results which restrict the block to approximate the input signal , then removes the portion of well-estimated signal from , therefore providing a hierarchical decomposition. is the partial forecast based on the decomposed input and the final forecast is the sum of all partial forecasts.
Subsequently, N-HiTs [57] redefine the N-BEATS by incorporating subsampling layers before the fully connected blocks, which enhances the input decomposition via multi-frequency data sampling and future predictor via multi-scale interpolation. DEPTS [58] puts forward a novel decoupled formulation for periodic time series by introducing the periodic state as a hidden variable and then develops a deep expansion module on top of residual learning to conduct layer-by-layer expansions between observed signals and hidden periodic states. Similarly, DEWP [59] is a also stack-by-stack expansion model to handle multivariate time series data, where each stack consists of a variable expansion block to capture dependencies among multiple variables and a time expansion block to learn temporal dependencies.
3.2.3 Matrix Factorization
The above-mentioned two decomposition methods are proposed for univariate series or applied to multivariate series in a variate-independent way. Here, we discuss a factorization-based decomposition for multivariate series. Specifically, many multivariate time series data in real-world scenarios can also be referred to as high-dimensional data. They can be formalized in the form of a matrix, whose rows correspond to variate and columns correspond to time points. Since variables in multivariate time series tend to be highly correlated, it can be possibly reduced to a more compact space. Matrix factorization methods [60] work by decomposing the high-dimensional series data into the product of two matrices in a lower-dimensional latent space. For a multivariate time series , as shown in Figure 2 the matrix can be approximated by the multiplications of two lower rank embedding matrix, , in which , and is a hyperparameter.
Besides the estimation, there are regularizers to avoid overfitting problems in factorization. Going beyond the canonical design that takes the squared Frobenius norm as regularizers, Temporal regularized matrix factorization (TRMF) [61] designs an autoregressive-based temporal regularizer to describe temporal dependencies among latent temporal embeddings. Further, [62] extended TRMF with a new spatial autoregressive regularizer to estimate low-rank latent factors by simultaneously learning the spatial and temporal autocorrelations. NoTMF [63] integrates the vector autoregressive process with differencing operations into the classical low-rank matrix factorization framework to better model real-world time series data with trend and seasonality. Eliminating the need for tuning regularization parameters, BTF [64] is a fully Bayesian model that integrates the probabilistic matrix factorization and vector autoregressive process into a single probabilistic graphical model. Instead of using an autoregressive-based temporal regularization, DeepGLO [65] utilizes a temporal convolution network for regularization to capture non-linear dependencies. LSTM-GL-ReMF [66] contains an LSTM-based temporal regularizer to learn complex long-term and short-term non-linear temporal correlations and a Graph Laplacian spatial regularizer [67] to capture spatial correlations.
3.3 Fourier Analysis
Fourier analysis [68, 69] can convert a physical signal into the Fourier domain to highlight the inherent frequency properties of the original data and has been a well-acknowledged analysis tool in extensive areas. Since time series are usually recorded as a sequence of discrete time points by sampling the original continuous signals, Fourier analysis has become one of the mainstream tools in time series modeling and has been demonstrated favorable effectiveness and efficiency [70, 71]. Introducing the Fourier domain not only augments the representation of the original series but also provides a global view since the frequency spectrum distribution, which can indicate essential periodic properties of time series. In practice, Fast Fourier Transform (FFT) [72] and Wavelet Transform (WT) [73] as the basic algorithms connecting the discrete temporal domain to the frequency domain, have gained increasing popularity in the modular design of deep time series models [74, 75, 76, 77, 78, 79, 80, 81]. Existing approaches can be roughly divided into two categories: time-domain and frequency-domain modeling.
3.3.1 Time-Domain Modeling
The fundamental principle behind the Fourier transform is that sequential data can be decomposed and represented by a series of periodic signals. Consequently, it can be used to identify potentially dominant periods and their corresponding frequencies in the data by analyzing the highest amplitude components. As a typical practice, TimesNet [3] employs the Fast Fourier Transform (FFT) to extract the most significant frequencies with the highest amplitude values, subsequently reshaping the 1D time series data into a 2D space based on the identified periods for better representation learning. Following TimesNet, PDF [82] posits that frequencies with larger values facilitate a more discernible distinction between long-term and short-term relationships.
In addition to exploiting the information of the sequence obtained by the Fourier Transformer, some works attempt to perform efficient computation through the Fast Fourier Transformer. Auto-correlation is a fundamental concept in time series analysis that measures the dependence between observations at different time points within a sequence of data. The Wiener-Khinchin theorem [83] provides a mathematical relationship between the auto-correlation function and the power spectral density (PSD) of a stationary random process, where the auto-correlation function represents the inverse Fourier transform of the PSD. Taking the data as a real discrete-time process, Autoformer [22] proposes an Auto-Correlation mechanism with an efficient Fast Fourier Transforms to capture the series-wise correlation.
The frequency-domain representation provides information about the amplitudes and phases, where low-frequency components correspond to slower variations or trends in the signal, and high-frequency components capture fine details or rapid variations. A significant body of work has focused on leveraging frequency-domain information to enhance the model’s capability in capturing temporal dependencies. FiLM [84] introduces Frequency Enhanced Layers (FEL) which combine Fourier analysis with low-rank approximation to keep the part of the representation related to low-frequency Fourier components and the top eigenspace to effectively reduce the noise and boost the training speed. FITS [85] integrates a low-pass filter (LPF) to eliminate high-frequency components above a specified cutoff frequency, thereby compressing the model size while preserving essential information. From an opposite idea, FEDformer [86] posits that retaining only low-frequency components is insufficient for time series modeling, as it may dismiss important fluctuations in the data. Based on the above considerations, to capture the global view of time series, FEDformer represents the series by randomly selecting a constant number of Fourier components, including both high-frequency and low-frequency components.
3.3.2 Frequency-Domain Modeling
Building on time-frequency analysis in signal processing, several approaches have been developed to study time series simultaneously in both the time and frequency domains. ATFN [87] comprises an augmented sequence-to-sequence model that learns the trending features of complex non-stationary time series, along with a frequency-domain block designed to capture dynamic and intricate periodic patterns. TFAD [88] introduces a time-frequency analysis-based model that employs temporal convolutional networks to learn both time-domain and frequency-domain representations.
Some works have developed specialized deep learning architecture to process the frequency domain of time series. STFNet [75] applies Short-Time Fourier Transform to input signals and applies filtering, convolution, and pooling operations directly in the frequency domain. StemGNN [28] combines Graph Fourier Transform (GFT) and Discrete Fourier Transform to model both inter-series correlations and temporal dependencies. EV-FGN [89] uses a 2D discrete Fourier transform on the spatial-temporal plane of the embeddings and performs graph convolutions for capturing the spatial-temporal dependencies simultaneously in the frequency domain. FreTS [90] leverages Discrete Fourier Transform (DFT) to transform the data into the frequency domain spectrum and introduces frequency domain MLPs designed for complex numbers with separated modeling for the real parts and the imaginary parts. FCVAE [91] integrates both the global and local frequency features into the condition of Conditional Variational Autoencoder (CVAE) concurrently. Recent TSLANet [92] propose a lightweight Adaptive Spectral Block (ASB) to replace the self-attention mechanism, which is achieved via Fourier-based multiplications by global and local filters. FourierDiffusion [93] explores extending the score-based SDE formulation of diffusion to complex-valued data and therefore implements time series diffusion in the frequency domain.
4 Model Architectures
As we have discussed in Section 2, the time series model needs to unearth the intrinsic temporal dependencies and variate correlations that lie in observations. In this section, we provide a technical review of the existing deep time series models. As we have presented in Figure 3, existing works can be classified into five categories based on their backbone architecture, namely MLP-based, RNN-based, CNN-based, GNN-based, and Transformer-based.
4.1 Multi-Layer Perceptrons
As a representation of traditional statistical time series models, the Auto-regressive (AR) model assumes that the model output depends linearly on its own historical values. Inspired by the remarkable performance of auto-regressive models, Multi-Layer Perceptrons (MLP) have become a popular architecture for modeling time series data.
As a representative work of linear-based models, N-BEATS [24] is a pure MLP-based deep time series model without any time-series-specific knowledge to capture the temporal patterns in time series. Specifically, as described in Equ. (4) N-BEATS consists of deep stacks of fully-connected layers with two residual branches in each layer, one is for the backcast prediction and the other one is the forecast branch. Extending the idea of neural basis expansion analysis, N-HiTs [57] use multi-rate signal sampling and hierarchical interpolation and N-BEATSx [94] incorporate exogenous variables to enhance the prediction.
Recent research by DLinear [52], also referred to as LTSF-Linear, challenges the effectiveness of complicated deep architecture in temporal modeling. It argues a simple linear regression in the raw space that achieves remarkable performance in both modeling and efficiency. As illustrated in Figure 4, prevalent MLP-based deep time series models consist of simple linear layers primarily designed for forecasting tasks. Also lightweight but effective, FITS [85] advocates time series analysis can be treated as interpolation exercises within the complex frequency domain and further introduces a complex-valued linear layer to learn amplitude scaling and phase shift in the frequency domain. Inspired by MLP-Mixer [95] in computer vision, several works have attempted to utilize MLPs to model both temporal and variate dependencies. TSMixer [96] contains interleaving time-mixing and feature-mixing MLPs to extract information from different perspectives. To better model the global dependencies in time series data, FreTS [90] investigates the learned patterns of frequency-domain MLPs which are operated on both inter-series and intra-series scales to capture channel-wise and time-wise dependencies in multivariate data.
Recent works have moved beyond using simple linear layers over discrete time points. TimeMixer suggests that time series exhibit distinct patterns in different sampling scales and proposes an MLP-based multiscale mixing architecture. TiDE [97] incorporates exogenous variables to enhance the time series prediction. Based on Koopman theory and Dynamic Mode Decomposition (DMD) [98], which is a dominant approach for analyzing complicated dynamical systems, Koopa [99] hierarchically disentangles dynamics through an end-to-end predictive training framework and can utilize real-time incoming series for online development.
4.2 Recurrent Neural Networks
Recurrent Neural Networks (RNNs) are specifically designed to model sequential data [100, 101, 102], such as natural language processing [103] and audio modeling [104]. Since time series are also serial in nature, RNNs have emerged as a popular choice for analyzing time series data [105]. Existing RNN-based deep time series models focus on combating the gradient vanishing problem caused by the vanilla recurrent structure and modeling the mutual correlation among multivariate variables. Previous works [106, 107, 108, 109, 110] use variants of RNN to model temporal dependencies.
LSTNet [111] combines the recurrent structure with the convolutional layer to capture both the short-term local dependency between variables and long-term patterns for time series. Moreover, a novel recurrent-skip component based on the periodic pattern is introduced to alleviate gradient vanishing in modeling long-term dependencies. Similarly, DA-RNN [112] combines the recurrent unit with a dual-stage attention mechanism to adaptively extract relevant series at each time step. Beyond deterministic forecasts, DeepAR [113] proposes an auto-regressive recurrent network model to predict the probability distribution of further time points. Technologically, it learns not only the seasonal behavior with time series but dependencies on given covariates across time series, allowing the model to make predictions even when there is little or no historical data.
Also based on Markovian state representation, the State Space Model (SSM) [114] is another classical mathematical framework that captures the probabilistic dependence between observed measurements in stochastic dynamical systems. Concretely, a single-input single-output (SISO) linear state space model is defined as follows:
where are input signal, latent state, and output signal respectively. The system is characterized by the matrices can be learned by the deep neural network.
SSMs have proven their effectiveness and efficiency in processing well-structured time series data, but traditional approaches have to refit each time series sample separately and therefore cannot infer shared patterns from a dataset of similar time series. With the rise of deep learning models, modern SSMs are often implemented in a recurrent manner. By adapting and propagating a deterministic hidden state, RNNs are able to represent long-term dependencies in continuous data which offer an alternative to classical state space models. Therefore, some work [115, 116] have attempted to fuse classical state space models with deep neural networks. Representative like Deep State Spaces Model (DSSM) [117], using a recurrent neural network (RNN) to parametrize a particular linear SSM, takes advantage of incorporating structural assumptions and learning complex patterns. Structured State Space sequence model (S4) [118] introduces a new parameterization for the SSM by conditioning matrix with a low-rank correction, allowing it to be diagonalized stably, which empowers the model with better long-term modeling capacity. Similar to S4, LS4 [119] is a generative model with latent space evolution following a state space ordinary differential equations (ODE).
Recent work on Mamba [120] has emerged as a powerful method for modeling long-context sequential data while scaling linearly with sequence length. Utilizing a simple selection mechanism that parameterizes the SSM parameters based on the input, Mamba can discern the importance of information in a manner similar to the attention mechanism, posing a potentially effective way to sequential modeling.
4.3 Convolutional Neural Networks
Since the semantic information of time series is mainly hidden in the temporal variation, Convolutional neural networks (CNN) [18, 121] have become a competitive backbone for their ability to capture local features and pattern recognition. By leveraging convolutions and hierarchical feature extraction, CNNs have shown remarkable success in various computer vision tasks, such as image classification [122], segmentation [123] and object detection [124].
Considering the temporal continuity of time series data, previous works [125, 126, 127] apply one-dimensional CNN (1D CNN) to capture the local patterns of time series data. Recent SCINet[128] applies normal convolutions with a hierarchical downsample-convolve-interact architecture to capture dynamic temporal dependencies at different temporal resolutions of time series data. Inspired by the idea of masked convolution [129], Wavenet[130] introduces causal convolution and dilated causal convolution to model long-range temporal causality. Similar to Wavenet, Temporal Convolutional Networks (TCN) [131] uses a stack of dilated convolutional kernels with progressively enlarged dilation factors to achieve a large receptive field. However, the limited receptive field of TCN makes it difficult for them to capture global relationships in time series data. Based on TCN, MICN[132] is a local-global convolution network that combines different convolution kernels to model temporal correlation from a local and global perspective. ModernTCN [133] boosts the traditional TCN to capture cross-time and cross-variable dependency by DWConv and ConvFFN separately. Considering that DWConv is proposed to learn temporal information, it is operated variate-independently to learn the temporal dependency of each univariate time series.
Beyond 1D space, motivated by the periodicity properties of time series data, TimesNet [3] transforms the 1D time series data into a set of 2D tensors in each TimesBlock based on the estimated period lengths, where the inter-period variations are presented in tensor columns and inner-period ones are shown in tensor rows. Here is a hyperparameter, corresponding to multiple 1D-to-2D transformations with different periods. Then it applies inception block [134, 135] to process the transformed 2D tensors, which can be summarized as:
where is the -th transformed 2D tensor. After passing through the inception block , the learned 2D representations are transformed back to 1D for aggregation. These transformations enable TimesNet to effectively capture both multi-scale intraperiod-variation and interperiod-variation simultaneously. Furthermore, by leveraging hierarchical convolutional layers, TimesNet is capable of learning both high-level and low-level representations, facilitating comprehensive time series analysis across four distinct tasks.
4.4 Graph Neural Networks
Analyzing multivariate time series data is often challenging due to the complex and often non-linear correlations between variables. To address this challenge, Graph neural networks (GNNs) [136, 137] have been widely adopted in time series analysis. By modeling multivariate data as a spatiotemporal graph, where each node represents a variable, GNNs can extract relationships among neighboring nodes and capture the temporal evolution of node attributes over time, thereby providing a robust framework for understanding the underlying dynamics of multivariate time series.
The core goal of GNN architecture is to model the underlying topological relations in multivariate data, therefore existing GNN-based works can be roughly divided into two categories based on whether graph structure is part of the input into the model. DCRNN [138] models the spatial dependency of traffic as a diffusion process on a directed graph and uses diffusion convolution to capture the spatial dependency, alongside a recurrent neural network to capture temporal dynamics. Similarly, STGCN [139] integrates graph convolutional networks to model the spatial dependencies among traffic sensors with temporal convolutions to capture the temporal dependencies in the traffic time series data. Graph WaveNet [140] combines graph convolution with dilated casual convolution and learns an adaptive dependency matrix through node embedding, enabling the model to automatically capture hidden spatial dependencies in spatial-temporal graph data. Similarly, AGCRN [141] enhances the traditional graph convolutional network with node adaptive parameter learning and data-adaptive graph generation modules, allowing for the automatic capture of spatial and temporal correlations without a pre-defined graph structure. MTGNN [142] introduces a graph learning layer to adaptively learn the graph adjacency matrix, thereby capturing hidden relationships among multivariate time series data. STFGNN [143] employs a Spatial-Temporal Fusion Graph Neural Network with a generated temporal graph to learn localized spatial-temporal heterogeneity and global spatial-temporal homogeneity. StemGNN [28] leverages the advantages of both the Graph Fourier Transform (GFT) and the Discrete Fourier Transform (DFT), modeling multivariate time series in the spectral domain.
4.5 Transformers
In the view of the great success in the field of natural language processing [16, 144, 145, 146, 147] and computer vision [19, 148, 149, 150], Transformers have also emerged as a powerful backbone for time series analysis. Benefiting from the self-attention mechanism [151], Transformer-based models can capture long-term temporal dependencies and complex multivariate correlations. As overviewed in Figure 8, existing Transformer-based time series models can be categorized based on the granularity of representation used in the attention mechanism, namely point-wise, patch-wise, and series-wise approaches.
| Category | Method | Architecture | Embeddding | Attention Mechanism |
| Vanilla | Transformer | Enc-Dec | Standard | |
| Point-wise | LogSparse [152] | Dec-only | Standard | |
| Informer [21] | Enc-Dec | Standard | ||
| Pyraformer [153] | Enc-Dec | Standard | ||
| Patch-wise | Autoformer [22] | Enc-Dec | Standard | |
| Crossformer [29] | Enc-Dec | Patch-Wise | ||
| PatchTST [23] | Enc-only | Patch-Wise | ||
| Variate-wise | iTransformer [30] | Enc-only | Variate-Wise |
4.5.1 Point-wise Dependency
Due to the serial nature of time series, most existing Transformer-based works use a point-wise representation of time series data and apply attention mechanisms to capture the correlations among different time points. Among these point-wise modeling approaches, Data Embedding is a crucial component that maps the value of time series data to a high-dimensional representation. Given time series with corresponding time stamp information , where is the variate number and is the types of time stamps, the embedding module can be summarized as follow:
where and is the dimension of the emebedded representation, value projection and timestamp embedding are implemented by channel-dimension linear layers, and denotes the absolute position embedding to preserve the sequential context of input series.
To better apply the Transformer architecture to the time series domain, researchers have explored two aspects: designing pre-processing modules and modifying the attention mechanisms. As we have discussed in Section 3.1, past RevIN [40] and Stationary [41] achieved superior performance by introducing Normalization and De-Normalization modules before and after the vanilla Transformer. Besides, Stationary [41] further proposes De-stationary Attention to avoid the over-stationarization problem.
Given that the canonical attention approach leads to a quadratic computational complexity, numerous efficient Transformers [21, 22, 153, 86] have been proposed to mitigate the complexity caused by point-wise modeling, which is summarized in Table II. LogSparse [152] proposes Convolutional Self-Attention to replace canonical attention by employing causal convolutions to produce queries and keys in the self-attention layer. Informer [21] introduces a Query Sparsity Measurement, where a larger value indicates a higher chance to contain the dominant information in self-attention. Based on the proposed sparsity measurement, it further designs a ProbSparse self-attention only using top queries with the biggest measurement results, which can reduce the quadratic complexity in time and memory. Pyraformer [153] constructs a multi-resolution -ary tree and develops a Pyramidal Attention Mechanism, in which every node can only attend to its neighboring, adjacent, and children nodes. With the calculated mask for attention, Pyraformer can capture both short- and long-temporal dependencies with linear time and space complexity.
4.5.2 Patch-wise Dependency
Patch-based architectures play a crucial role in the Transformer models for both Natural Language Processing (NLP) [16] and Computer Vision (CV) [19]. Since point-wise representations are insufficient to capture local semantic information in temporal data, several studies [23, 154, 155] have been devoted to exploring patch-level temporal dependencies within time series data.
Pioneer work Autoformer [22] proposes an Auto-Correlation Mechanism, which captures the series-wise dependencies of time series to replace canonical point-wise self-attention. Based on the stochastic process theory [156], Auto-Correlation utilizes the Fast Fourier Transform to discover the time-delay similarities between different sub-series. A time delay module is further proposed to aggregate the similar sub-series from underlying periods instead of the relation between scattered points, which firstly explores the sub-series level modeling in Transformer-based models.
Different from modifying the attention mechanism, most recent works utilize patch-wise representations of time series data and perform a self-attention mechanism to capture patch-wise dependencies [23, 54, 157]. PatchTST [23] and follow-up works split time series into a sequence of overlapped patches and embed each patch following:
Assume , is patch length and the corresponding number of patches split, and denotes the -th patch with sequence length . The patches are mapped to the latent space through a temporal linear projection and a learnable position embedding .
Based on the vanilla attention mechanism, PatchTST [23] learns the patch-wise dependencies. Going beyond PatchTST, recent Pathformer [158] proposes a multi-scale Transformer-based model with adaptive pathways. Based on the patch division of different scales, the adaptive pathways select the patch sizes with the top weights generated by the router to capture multi-scale characteristics.
The success of PatchTST also benefits from channel-independence design, where each temporal patch-level token only contains information from a single series. In addition to capturing the patch-level temporal dependencies within one single series, recent approaches [54, 157] have endeavored to capture interdependencies among patches from different variables over time. Crossformer [29] introduces a Two-Stage Attention layer containing a Cross-Time Stage and a Cross-Dimension Stage to efficiently capture the cross-time and cross-variate dependencies between each patch token. For the obtained embedded vector , the overall attention stage can be described as follow:
where is a learnable vector array used as a router to gather information from all dimensions and then distribute the gathered information.
4.5.3 Series-wise Dependency
Further expanding the receptive field, there are also some works that attempt to use the tokenization of the whole time series to capture inter-series dependencies. iTransformer [30] introduce to multivariate data, and for -th variable , it can be simply formulated as follows:
where is instantiated as trainable linear projector. Based on the global representations of each series, iTransformer utilizes the vanilla Transformer without any architectural modifications to capture mutual correlations in multivariate time series data. Similarly, TimeXer[159] focuses on forecasting with exogenous variables and utilizes patch-level and series-level representations for endogenous and exogenous variables, respectively. Additionally, an endogenous global token is introduced to TimeXer, which serves as a bridge in-between and therefore captures intra-endogenous temporal dependencies and exogenous-to-endogenous correlations jointly.
5 Time Series Library
Time series analysis has emerged as an important research area, attracting significant attention from both academia and industry. Recently, extensive exploration of deep learning based methods for time series analysis has resulted in significant advances. However, the issue of fair benchmarking poses a pressing challenge in this domain. The absence of fair, rational, and comprehensive benchmarks can lead to biased comparisons between different methods and hinder accurate evaluation of their effectiveness, potentially inflating domain advances or hindering practical applications. This presents a substantial obstacle to understanding advances and fostering robust development within the field.
In the domain of time series analysis, several benchmarks have been proposed, such as DGCRN [160], LibCity [161], DL-Traff [162], TS-bench [163], and BasicTS [14]. More specifically, Autoformer [22] proposed a standard long-term forecasting benchmark covering different practical applications. Further, to verify the generality of different time series analysis models, TimesNet [3] builds a more comprehensive model generalization benchmark covering five mainstream time series analysis tasks. However, these benchmarks typically have some limitations. One issue with current time series benchmarks is their limited coverage of time series analysis tasks and specific domains, which limits their practical applications. Moreover, these benchmarks often fail to provide detailed discussions and comprehensive summaries of task types, model architectures, and specific baseline methods. As a result, they do not effectively guide the design of more efficient time series analysis methods or drive further development in the field.
To effectively address these issues, we introduce and implement Time Series Library (TSLib), a benchmark for fair and comprehensive comparing and evaluating the performance of deep time series models across various time series analysis tasks. As shown in Figure 9, TSLib encompasses a unified model experiment pipeline, standardized evaluation protocols, extensive and diverse real-world datasets, mainstream and advanced time series analysis models, and unified experimental validation and analysis process.
In our Time Series Library, we meticulously followed the official codes and implemented 24 widely used and advanced deep time series analysis models. These models are derived from four canonical deep learning architectures. Users can choose from these models based on their specific practical usage scenarios. The code is available at https://github.com/thuml/Time-Series-Library.
As follows, we will provide a detailed description of our TSLib, including the design and implementation principles (Section 5.1), the evaluation protocols and metrics (Section 5.2), the dataset descriptions (Section 5.3), and the main results of models with different architectures (Section 5.4).
5.1 Design and Implementation Principle
TSLib is designed based on the well-established factory pattern and implements a unified interface between data and model objects, thus enabling a clear separation between deep model creation and usage, promoting modularity and flexibility. By loading different data and model objects and combining specific task heads during model training, TSLib enables different datasets and models to be shared and extended, allowing easy switching between various time series analysis tasks. These design and implementation principles provide enhanced flexibility and scalability for our TSLib. Furthermore, as illustrated in Figure 9, TSLib introduces a unified experimental pipeline covering the overall process of the model training and evaluation, which includes data source, data processing, model training and analysis, and model performance evaluation.
Data Source Our TSLib provides extensive support for a wide range of diverse and multi-type datasets in a variety of formats, including ”.csv”, ”.npz”, ”.txt”, etc. As shown in Figure 9 and Table III, TSLib currently supports more than 30 datasets with different sampled frequencies across four mainstream time series analysis tasks, all derived from real-world scenarios in domains such as energy, transportation, economics, weather, and medicine, etc. Moreover, TSLib excels in scalability, allowing for the effortless integration of new data sources of different data types.
Data Processing Data processing plays a pivotal role in guaranteeing stable training within the realm of time series analysis. Within the Time Series Library, a multitude of data processing steps are conducted, including time window splitting, data batch generation, etc. Subsequently, the raw data is partitioned into separate sets for training, validation, and testing purposes, enabling streamlined model training and equitable comparisons. These steps serve as indispensable prerequisites for attaining precise and dependable results across a range of diverse time series analysis tasks.
Moreover, our TSLib provides additional support for numerous crucial and effective data processing strategies based on different model design principles [39], [40], [41], [45] to enhance model performance and training efficiency. We encapsulate these various design strategies within our basic data processing layer, encompassing techniques such as data normalization, time-frequency decomposition, Fourier analysis, and more. When utilizing TSLib, users have the flexibility to select these strategies to improve training effect based on their specific requirements and objectives.
Model Training and Analysis After the data processing phase, the raw time series data is transformed into the desired format for model training. Model training forms the crux of the entire experiment pipeline, where we fine-tune the model parameters based on the input to predict the output with minimal error. Our primary goal during model training is to obtain the best possible trainable parameters that result in a significant improvement in model performance. Each model has its own unique design and training objective. The model analysis procedure is to determine the optimal model parameters by comparing the correlation between the training and validation losses. Our TSLib includes complete log printing and result storage functions record and evaluate the training process. By employing rational model analysis techniques, we can efficiently obtain models with superior performance and stronger generalization.
Performance Evaluation Model evaluation is a crucial step in verifying the effectiveness and generalization of trained time series models. It involves model prediction and performance evaluation, providing insights into the efficacy of the trained model. In TSLib, we provide evaluation support for four mainstream time series analysis tasks: classification, imputation, forecasting (long-term or short-term), and anomaly detection. Each task comes with its specific evaluation metric, enabling a comprehensive assessment of the performance of models. These metrics play a crucial role in determining the effectiveness of the trained model and its suitability for the intended task.
| Task | Dataset | Dimension | Length | Domain | Size |
| Classification | EthanolConcentration | 3 | 1,751 | Alcohol Industry | 20.3 MB |
| FaceDetection | 144 | 62 | Face (250 Hz) | 789.1 MB | |
| Handwriting | 3 | 152 | Motion | 3.9 MB | |
| Heartbeat | 61 | 405 | Health (0.061 secs) | 87.8 MB | |
| JapaneseVowels | 12 | 29 | Voice | 1.1 MB | |
| PEMS-SF | 963 | 144 | Transportation (1 day) | 420.1 MB | |
| SelfRegulationSCP1 | 6 | 896 | Health (256 Hz) | 17.8 MB | |
| SelfRegulationSCP2 | 7 | 1,152 | Health (256 Hz) | 17.7 MB | |
| SpokenArabicDigits | 13 | 93 | Voice (11025 Hz) | 37.6 MB | |
| UWaveGestureLibrary | 3 | 315 | Gesture | 3.4 MB | |
| Imputation | ETTh1, ETTh2 | 7 | 17,420 | Electricity (1 hour) | 10.4 MB |
| ETTm1, ETTm2 | 7 | 69,680 | Electricity (15 mins) | 2.6 MB | |
| Electricity | 321 | 26,304 | Electricity (1 hour) | 95.6 MB | |
| Weather | 21 | 52,696 | Environment (10 mins) | 7.2 MB | |
| Long-term Forecasting | ETTh1, ETTh2 | 7 | 17,420 | Electricity (1 hour) | 10.4 MB |
| ETTm1, ETTm2 | 7 | 69,680 | Electricity (15 mins) | 2.6 MB | |
| Electricity | 321 | 26,304 | Electricity (1 hour) | 95.6 MB | |
| Weather | 21 | 52,696 | Environment (10 mins) | 7.2 MB | |
| Traffic | 862 | 17,544 | Transportation (1 hour) | 136.5 MB | |
| Exchange | 8 | 7,588 | Economic (1 day) | 623 KB | |
| ILI | 7 | 966 | Health (1 week) | 66 KB | |
| Short-term Forecasting | M4-Yearly | 1 | 6 | Demographic | 589.5 MB |
| M4-Quarterly | 1 | 8 | Finance | ||
| M4-Monthly | 1 | 18 | Industry | ||
| M4-Weakly | 1 | 13 | Macro | ||
| M4-Daily | 1 | 14 | Micro | ||
| M4-Hourly | 1 | 48 | Other | ||
| Anomaly Detection | SMD | 38 | 100 | Industry (1 min) | 436.4 MB |
| MSL | 55 | 100 | Industry (1 min) | 58.2 MB | |
| SMAP | 25 | 100 | Industry (1 min) | 113.0 MB | |
| SwaT | 51 | 100 | Industry (1 min) | 903.2 MB | |
| PSM | 25 | 100 | Industry (1 min) | 107.1 MB |
5.2 Evaluation Protocols
In order to conduct a fair and comprehensive model performance verification, our Time Series Library is designed to provide standardized evaluation protocols for four mainstream time series analysis tasks following [3]. The primary goal of these standardized and unified evaluation protocols is to quantify the effectiveness of different time series analysis methods with varying architectures. Additionally, they provide valuable insights into the strengths and limitations of different methods across diverse time series analysis tasks. By establishing these standardized evaluation protocols, we aim to promote fair comparisons between different methods and improve our understanding of their performance in various time series analysis scenarios.
For long-term forecasting and imputations, we rely on Mean Square Error (MSE) and Mean Absolute Error (MAE) as the primary evaluation metrics. These metrics help us accurately assess the accuracy of our predictions and imputations. For short-term forecasting, we use the Symmetric Mean Absolute Percentage Error (SMAPE) and Mean Absolute Scaled Error (MASE) as metrics, which focus on absolute errors and reduce the impact of outliers, providing reliable evaluations of forecast accuracy across different datasets and methodologies. In the case of time series classification tasks, we utilize Accuracy as the evaluation metric. Accuracy measures the overall prediction performance by calculating the ratio of correctly classified samples to the total number of samples. For anomaly detection, we employ the F1 score to validate the identification of abnormal values. The F1 score represents a balanced combination of precision and recall, offering a comprehensive assessment of a classifier’s performance, especially when dealing with imbalanced classes in the context of anomaly detection.
5.3 Datasets
TSLib includes a variety of mainstream datasets across different numbers of samples and categories, a richness of tasks, and a diversity of domains. In this section, we will focus on introducing representative datasets from various time series analysis tasks included in the Time Series Library.
Classification Time series classification aims to assign a label or category to a time series based on its temporal features. To evaluate this capability, we selected ten multivariate datasets from the UEA Time Series Classification Archive [164], supported in our TSLib. These datasets cover a range of practical tasks, including gesture, action, audio recognition, and medical diagnosis through heartbeat monitoring. We pre-processed the datasets according to the descriptions provided in [165]. Detailed dataset descriptions are shown in Table III.
Imputation Due to glitches, the collected time series data may contain partially missing values, posing a challenge for time series analysis. Therefore, time series imputation is a crucial task in real-world applications, aiming to fill in missing values within a time series based on contextual observations from the data. For our benchmark, we selected Electricity Transformer Temperature (ETT) [21], Electricity, and Weather to evaluate the performance of time series imputation tasks with different missing ratios.
Forecasting Time series forecasting is an essential task in time series analysis, and it has been widely explored in academic and industry domains. By leveraging historical patterns and trends, the model can predict future values or trends of a time series. Time series forecasting can be broadly divided into two types: long-term forecasting and short-term forecasting. For the long-term time series forecasting task, a wide range of datasets are included in our benchmark, including Electricity Transformer Temperature (ETT), Electricity, Weather, Traffic, Exchange, and Illness (ILI). For the short-term forecasting task, we selected the M4 dataset [166], which comprises six sub-datasets with varying sampling frequencies and domains.
Anomaly Detection Anomaly detection involves identifying unusual or abnormal patterns in a time series. These anomalies can indicate critical events, faults, or outliers that require attention or further investigation. There are some mainstream anomaly detection datasets supported in TSLib, such as Server Machine Dataset (SMD) [167], Mars Science Laboratory rover (MSL) [168], Soil Moisture Active Passive satellite (SMAP) [168], Secure Water Treatment (SWaT) [169], and Pooled Server Metrics (PSM) [170] which are collected from a variety of industrial scenarios.
5.4 Main Results
To examine the strengths and limitations of various methods in mainstream time series analysis tasks, we select twelve representative models from our TSLib. These models encompass four popular deep model architectures and address tasks, including long-term and short-term forecasting, classification, imputation, and anomaly detection.
Baselines To conduct a fair comparative analysis and thoroughly explore the effectiveness of different model architectures in various time series analysis tasks, we conduct comparative comparison experiments using state-of-the-art models designed based on different deep architectures. As shown in Figure 11, we select several advanced and representative Transformer-based models: iTransformer [30], PatchTST [23], Autoformer [22], Non-Stationary Transformer (Stationary) [41], and FEDformer [86] to verify the performance. Additionally, we consider TimesNet [3] and SCINet [128] as the CNN-based models to compare. For RNN-based models, we included the novel and effective Mamba [120]. Finally, we include DLinear [52], N-BEATS [24], and TiDE [97] as representative MLP-based models for analysis. It is important to mention that TiDE [97] is designed to be dependent on specific timestamps, and cannot be easily adapted to part tasks without timestamps, such as short-term forecasting, anomaly detection, and classification.
Unified Experimental Settings For the long-term forecasting task, we conducted two experimental settings to ensure a fair and comprehensive comparison: unified hyperparameter and hyperparameter searching. For the unified hyperparameter, we conducted experiments using a range of standardized hyperparameters set across different datasets. This allows us to accurately evaluate the relative performance of time series models with different deep architectures while keeping other factors constant. As for the “hyperparameter searching” scenario, we conducted separate hyperparameter searches for different model architectures and time series datasets. This approach enabled us to identify the best performance of different time series analysis models. By employing above both settings, we obtain a comprehensive understanding of the forecasting performance of different time series models. For the remaining tasks, we maintained the standard experimental settings as outlined in [3] to validate the performance of different time series models.
Overall Results Based on the overall results of the models across different architectures in Figure 11, we surprisingly find that the MLP-based models, which are generally simpler and have lower computational overhead, perform well on the time series forecasting task. However, these models appear to be less effective in other types of tasks, which requires the model to learn more informative representations. On the contrary, the CNN-based models demonstrate more comprehensive capabilities and excel in classification, imputation, and anomaly detection tasks. The RNN-based models, while performing well on anomaly detection tasks, show limited effectiveness compared to other model architectures. In contrast, the Transformer-based models have demonstrated highly competitive performance across various time series analysis tasks. This can be attributed to the powerful data modeling capabilities inherent in the transformer architecture, which contribute to its overall and consistently superior performance across diverse time series analysis tasks. It further shows that Transformer-based models hold significant research value and application potential in the field of time series analysis and have emerged as a particularly promising option in the time series domain.
As illustrated in Figure 10, we have also included more detailed results and a top three performance leaderboard for four representative time series analysis tasks. These results clearly show that the Transformer-based models, namely iTransformer [30] and PatchTST [23], exhibit superior forecasting capabilities compared to other models for both long-term and short-term forecasting tasks. This further proves that it is of great significance and value to explore different modeling methods of temporal tokens in time series. Additionally, TimesNet [3] shows a more comprehensive and effective performance covering time series classification, imputation, and anomaly detection tasks. It has pioneered a milestone in the general time series analysis model.
We believe that TSLib can provide useful start code, valuable insights on model properties and model selection guidance for future research and real-world applications.
6 Future Directions
In this section, we provide a discussion on the promising directions for time series analysis.
6.1 Time Series Pre-training
A pretraining-finetuning learning paradigm is a two-stage approach commonly used in nature language processing (NLP) [171], [16], [172], [173], [173] and computer vision (CV) [174], [175], [176], [177]. Pre-training establishes the basis of the abilities of Large Models through unsupervised learning [178, 179]. Fine-tuning can improve the performance of the pre-trained model on a specific task or domain.
Due to the limited availability of labeled datasets, self-supervised pre-training [175] has garnered significant attention and has been extensively investigated in the domains of natural language modeling and computer vision. Self-supervised pre-training paradigm significantly reduces labeling expenses and benefits for diverse downstream tasks. Notably, recent research efforts have introduced several self-supervised pre-training methods tailored for time series data, which can be primarily classified into contrastive learning [180] and masked time series modeling [16, 181].
Contrastive learning refers to learning the representations of data by contrasting between similar and dissimilar pairs, where similar sample pairs are learned to be close to each other and dissimilar pairs are far apart [180, 182, 183]. Although SimCLR [182] has demonstrated remarkable success in the domain of computer vision, directly applying SimCLR to the field of time series data yields unsatisfactory results due to the insufficient modeling of temporal dependencies. CPC [184] introduced contrastive predictive coding, which utilizes model-predicted features as positive samples and randomly-sampled features as negative samples, to obtain time series representations that are advantageous for downstream tasks. TimeCLR [185] proposes a DTW data augmentation to generate phase shift and amplitude-change phenomena which can preserve time series structure and feature information. TS-TCC [186] employs efficient data augmentations designed for time-series data, and learns discriminative representations from the proposed Temporal Contrasting module and Contextual Contrasting module. TS2Vec [187] employs a hierarchical contrastive learning method and defines the contrastive loss from both instance-wise and patch-wise perspectives across different augmented context views, resulting in a robust contextual representation for each timestamp. Furthermore. LaST [188] takes the idea of variational inference theory [189] and proposes seasonal-trend representations learning and disentanglement mechanisms. CoST [190] proposes a contrastive learning framework to learn disentangled seasonal-trend representations for long sequence time series data. TF-C [191] develop frequency-based contrastive augmentation to leverage rich spectral information and explore time-frequency consistency.
Masked modeling is a reconstruction-based method, which can predict a masked token in a sequence based on the context unmasked part [192, 16, 181]. TST [193] follows the pre-training paradigm proposed in BERT [16] and proposes a pre-training framework for multivariate time series. Further, PatchTST [23] segments the time series data into multiple non-overlapping patches and proposes a patch-level masked modeling approach. SimMTM [194] relates masked modeling to manifold learning and presents a neighborhood aggregation design for reconstruction based on the similarities learned in series-wise representation space. HiMTM [195] proposes a novel hierarchical masked time series pre-training framework to effectively capture the multi-scale characteristics of time series data. TimeSiam [196] constructs an asymmetric masking reconstruction task to capture intrinsic temporal correlations between randomly sampled past and current subseries and learn internal time-dependent representations based on Siamese networks.
6.2 Large Time Series Models
With the advent of Large Language Models (LLMs), the utilization of large-scale models to tackle time series downstream tasks has gained significant attention as the direction of future research. Current researches present the following two possible roadmaps to large time series models.
6.2.1 Time Series Foundation Models
Recent advancements in deep learning, particularly with the emergence of foundation models (FMs), have demonstrated significant progress in natural language processing (NLP) and computer vision (CV) domains [197, 179, 198]. Different from prior deep models, foundation models are pre-trained on massive amounts of data, which enables them to have a wide range of general knowledge learned from diverse domains. Given their success in capturing contextual information and semantic understanding, it is promising to explore a generalized time series foundation model that can effectively learn complex temporal dependencies and capture the underlying dynamics inherent in time series data. Early attempts such as TimeGPT [199], Lag-LlaMa [200], and Timer [155] focus solely on univariate time series data. Nevertheless, in real-world forecasting scenarios, it is crucial to involve additional information that is related to the temporal variation of the target time series and must be taken into account, such as weather conditions or holidays. MOIRAI [201] tries to flatten multivariate time series into a single sequence containing all variate, but its generalization capabilities to other downstream analysis tasks are under-explored. In addition to modeling inter-series dependencies, modeling the relationship between time series and external factors in other can achieve a better understanding of time series data. These external factors may be in the form of other modalities, such as text data or calendar data, and thus multimodal learning is a future trend in the development of a multi-modal time series foundation model.
6.2.2 Adaptation of Large Language Models
Large Language Models (LLMs) have made significant strides in solving various natural language processing tasks. Exemplified by the success of models like GPTs [146, 17] and LLaMA [202, 203], LLMs have proven adept at generalizing to unseen tasks by simply following provided prompts. Therefore, it has become a promising future research direction to unleash the power of LLMs in the field of time series. Here are two paradigms for adapting LLMs.
Fine-Tuning Pre-trained Language Models Based on the similar sequential nature, fine-tuning pre-trained language models to equip them with time series analysis capabilities has become a promising research topic. When applying LLMs to time series, it is essential to tokenize the time series before feeding it to a pre-trained model. Thus, adapting an LLM for time series included two key components: time series tokenization and efficient fine-tuning methods for time series analysis tasks. LLM4TS [204] proposes a two-stage fine-tuning strategy, including the time-series alignment stage to align LLMs with the nuances of time series data, and the fine-tuning stage for downstream tasks. LLMTime [205] treats time series forecasting as next-token prediction in text and attempts to encode time series data as a string of numerical digits. Recent Chronos [206] introduces a pre-trained probabilistic time series model based on existing Transformer-based language model architectures. Technologically, Chronos tokenizes time series values using scaling and quantization into a fixed vocabulary and trains the model on these tokenized time series via the cross-entropy loss. Benefiting from the generative capability of LLMs, most of the existing research focuses on time series forecasting tasks. GPT4TS [207] propose a unified framework for diverse time series analysis tasks by using a pre-trained GPT-2 model and fine-tuning the positional embeddings and the parameters of the layer normalization for each analysis task.
Prompting Large Language Models Recent Large Language Models exhibit the strong abilities of in-context learning [146] and instruction following [208]. Therefore, the paradigm of leveraging natural language instructions or task examples to guide the model in addressing novel tasks has emerged as a groundbreaking approach [209, 210, 211], which has become a potential solution for time series analysis tasks [212]. Recent literature, such as PromptCast [212], UniTime [213], and TimeLLM [214] focus on investigating a prompt template to enable LLMs to perform the forecasting task. There are other works represented by Autotimes [55, 215], that attempt to design soft prompts for time series data. However, existing prompting approaches are tailored for forecasting, and how to empower LLMs to other time series tasks besides forecasting is relatively unexplored.
6.3 Practical Applications
6.3.1 Handling Extremely Long Series
Deep time series models have demonstrated remarkable performance across a wide range of downstream tasks, their applicability to longer time series data is often limited by scalability and high computational complexity. In industrial time series analysis, high-frequency sampling results in lengthy historical data, impeding the practical implementation of advanced deep models. Existing methods usually include patching techniques to enable them to handle long sequences, and when the input length becomes longer, the patch length can be increased accordingly to reduce the computational complexity. However, model performance is closely tied to patch length; hence, solely increasing patch size to reduce complexity may compromise capabilities. Therefore, addressing the limitations of deep models in handling longer time series could be a promising topic.
6.3.2 Utilizing Exogenous Variables
Since variations within the time series are often influenced by external factors, it is crucial to include exogenous variables in the analysis to gain a more comprehensive understanding of these factors. Exogenous variables, which are widely discussed in time series prediction tasks, are included in the model for uniform training in modeling time series data, without requiring separate analysis. However, in practical applications, different from multivariate time series analysis, the main variables and covariates usually occupy different positions. Given the crucial role played by exogenous variables in real-world applications, it is essential to explore a unified framework for modeling the relationships between the endogenous and exogenous variants, which allows for a more comprehensive understanding of interrelations and causality among different variants, leading to better and more reliable model performance, as well as interpretability.
6.3.3 Processing Heterogeneous Data
In the field of time series analysis, there is still an unexplored area related to the modeling of heterogeneous time series data. Heterogeneous time series data encompasses a wide range of diverse characteristics, such as varying sampling rates, irregularities, and different length scales. These diverse features make it challenging to develop models that can effectively capture the underlying patterns and relationships within the data. Moreover, the need for fixed-size inputs in current deep learning models limits their ability to handle the dynamic nature of heterogeneous time series data. Addressing these challenges requires innovative approaches that can adapt to the unique nature of each individual time series while still capturing the overarching patterns across multiple series. This may involve developing new techniques for feature extraction, incorporating domain knowledge into model design, or exploring alternative architectures that are better suited to handle variable-length inputs.
As researchers continue to explore this unexplored area in time series analysis, there is potential for significant advances in areas such as finance, healthcare, and environmental monitoring. By improving our ability to model and analyze heterogeneous time series data, we can gain deeper insights into complex systems and make more informed decisions based on predictive analytics. Overall, further research in this area holds great promise for advancing our understanding of temporal data dynamics and enhancing the capabilities of time series modeling in real-world applications.
7 Conclusion
In this survey, we provide a systematic review of deep models in time series analysis and introduce Time-Series Library (TSLib) as a fair benchmark for deep time series models across various analysis tasks. Compared with previous reviews that focus on a specific analysis task or model architecture, this paper provides a comprehensive survey and overview of existing deep models for time series analysis, ranging from forecasting, classification, imputation, and anomaly detection. We first present a detailed review of the universal modules that are widely used among time series models, including normalization, decomposition, and Fourier analysis. Next, we summarize existing deep time series models from the perspective of backbone architecture. Based on the review of existing literature, we introduce a practical open-source library, Time Series Library (TSLib), which has included representative deep time series models that can be a fair evaluation benchmark in the field of time series analysis. Finally, we discuss future research directions for deep time series models based on the recent development of the AI community and the practical application needs of time series analysis in real-world scenarios.

### v2

@@
 Abstract
-Time series, characterized by a sequence of data points arranged in a discrete-time order, are ubiquitous in real-world applications.
-Different from other modalities, time series present unique challenges due to their complex and dynamic nature, including the entanglement of nonlinear patterns and time-variant trends.
-Analyzing time series data is of great significance in real-world scenarios and has been widely studied over centuries.
-Recent years have witnessed remarkable breakthroughs in the time series community, with techniques shifting from traditional statistical methods to advanced deep learning models.
+Time series, characterized by a sequence of data points organized in a discrete-time order, are ubiquitous in real-world scenarios.
+Unlike other data modalities, time series present unique challenges due to their intricate and dynamic nature, including the entanglement of nonlinear patterns and time-variant trends.
+Analyzing such data is of great significance in practical applications and has been extensively studied for centuries.
+Recent years have witnessed remarkable breakthroughs in the time series community, with techniques shifting from traditional statistical methods to contemporary deep learning models.
 In this paper, we delve into the design of deep time series models across various analysis tasks and review the existing literature from two perspectives: basic modules and model architectures.
-Further, we develop and release Time Series Library (TSLib) as a fair benchmark of deep time series models for diverse analysis tasks, which implements 24 mainstream models, covers 30 datasets from different domains, and supports five prevalent analysis tasks.
-Based on TSLib, we thoroughly evaluate 12 advanced deep time series models on different tasks.
-Empirical results indicate that models with specific structures are well-suited for distinct analytical tasks, which offers insights for research and adoption of deep time series models.
-Code is available at https://github.com/thuml/Time-Series-Library.
+Further, we develop and release Time Series Library (TSLib) as a fair benchmark of deep time series models for diverse analysis tasks.
+TSLib implements 30 prominent models, covers 30 datasets from different domains, and supports five prevalent analysis tasks.
+Based on TSLib, we thoroughly evaluate 13 advanced deep time series models across diverse tasks.
+Empirical results indicate that models with specific structures are well-suited for distinct analytical tasks, providing insights for research and adoption of deep time series models.
+Code and datasets are available at https://github.com/thuml/Time-Series-Library.
 Index Terms:
@@
 Compared to image and text data, which have objectively prescribed syntax or intuitive patterns, the semantic information of time series data is primarily derived from the temporal variation [3].
-This presents significant challenges in understanding the data, such as identifying sequential dependencies, trends, seasonal patterns, and complicated dynamics.
-Consequently, analyzing time series data requires sophisticated methods to capture and utilize these complex temporal representations.
-Given the crucial role of time series data in real-world applications [4, 5, 6], time series analysis has been a longstanding research direction.
-Time series analysis encompasses the process of analyzing the temporal variation to understand time series data and make accurate predictions and informed decisions.
-One of the essential cornerstone of time series analysis is discovering the underlying patterns in time series data, which involves the intricate temporal dependencies and variate correlations inherent within the data.
-By capturing these complex dependencies, time series models can effectively reveal the underlying dynamics, and facilitate various downstream tasks, including forecasting, classification, imputation, and anomaly detection.
-Traditional time series methods, such as AutoRegressive Integrated Moving Average (ARIMA) [1], Exponential Smoothing, and Spectral Analysis [7], have long served as stalwart tools in time series analysis.
-These models, grounded in statistical methodologies, have been instrumental in discovering patterns, trends, and seasonality within temporal variations.
-However, their capabilities are hindered due to the inherent limitations of capturing complex nonlinear relationships and long-term dependencies present in real-world time series data.
-The rigid assumptions of linearity and stationarity that underpin traditional models constrain their adaptability to eventful and evolving data flows.
-| Survey | Analysis Task | Model Architecture | Benchmark | |||||||
-| Forecasting | Classification | Imputation | Anomaly | MLP | CNN | RNN | GNN | Transformer | ||
-| Fawaz et al. (2019) [8] | ✓ | ✓ | ✓ | ✓ | ||||||
-| Braei et al. (2020) [9] | ✓ | ✓ | ✓ | ✓ | ||||||
-| Torres et al. (2021) [10] | ✓ | ✓ | ✓ | ✓ | ||||||
-| García et al. (2021) [11] | ✓ | ✓ | ✓ | ✓ | ||||||
-| Wen et al. (2022) [12] | ✓ | ✓ | ✓ | ✓ | ||||||
-| Jin et al. (2023) [13] | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
-| Shao et al. (2023) [14] | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||
-| Qiu et al. (2024) [15] | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||
-| Our Survey | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
-Deep models have garnered significant attention and achieved remarkable performance across various domains, including natural language processing (NLP) [16], [17], computer vision (CV) [18], [19], and recommendation systems [20].
-In recent years, deep learning models [21, 22, 23, 24, 3] have demonstrated their capability to capture the intricate dependencies within time series data, making deep learning models a powerful tool for time series analysis over traditional statistical methods.
-More recently, Transformer models with attention mechanisms, originally developed for natural language processing tasks, have presented stunning power in processing large-scale data [17] and have also been adapted for learning time series data.
+This presents significant challenges in understanding the data, notably in identifying sequential dependencies, trends, seasonal patterns, and complicated dynamics.
+Consequently, analyzing time series data necessitates sophisticated methods to extract its meaningful temporal representations.
+Time series analysis has long stood as a pivotal research area, owing to its indispensable role in real-world applications [4, 5, 6].
+It involves scrutinizing temporal variations in sequential data to uncover underlying patterns, thereby facilitating accurate predictions and informed decision-making.
+A fundamental aspect of this process lies in identifying intricate temporal dependencies and variable correlations inherent in the data.
+By capturing these entangled dependencies, time series models reveal underlying dynamics and support a broad spectrum of downstream tasks, including forecasting, classification, imputation, and anomaly detection.
+Traditional approaches, such as Autoregressive Integrated Moving Average (ARIMA) [1], Exponential Smoothing, and Spectral Analysis [7], have long served as stalwart tools for analyzing time series data.
+Grounded in statistical methodologies, these methods have been instrumental in uncovering patterns, trends, and seasonality within temporal variations.
+However, their rigid assumptions of linearity and stationarity limit their application to dynamic and evolving data flows, making them inadequate for capturing the intricate nonlinear relationships and long-term dependencies present in real-world time series data.
+| Survey |
+Analysis Task |
+Model Architecture |
+Benchmark | |||||||
+Forecasting |
+Classification |
+Imputation |
+Anomaly Detection |
+MLP |
+CNN |
+RNN |
+GNN |
+Transformer |
+Fawaz et al. (2019) [8] |
+Braei et al. (2020) [9] |
+Torres et al. (2021) |
+García et al. (2021) |
+Wen et al. (2022) [12] |
+Jin et al. (2023) [13] |
+Shao et al. (2023) [14] |
+Qiu et al. (2024) [15] |
+This Survey |
+Deep models have garnered wide attention and achieved remarkable performance across various domains, including natural language processing (NLP) [16], [17], computer vision (CV) [18], [19], and recommendation systems [20].
+In recent years, deep learning models [21, 22, 23, 24, 3] have proven highly effective in capturing the intricate dependencies within time series data, making them a powerful tool for time series analysis over traditional statistical methods.
+More recently, Transformer models with attention mechanisms, originally developed for natural language processing tasks, have exhibited unprecedented capability in processing large-scale data [17] and have also been adapted for learning time series data.
 These architectures offer the advantage of selectively focusing on different parts of the input sequence, allowing for more nuanced discovery of both temporal and variable dependencies in time series.
-Related Surveys Although various time series models designed for different analysis tasks have emerged in recent years, there is a lack of a comprehensive overview of existing methods, covering both tasks and models.
+Although various time series models designed for different analysis tasks have emerged in recent years, there is a lack of a comprehensive overview of existing methods, covering both tasks and models.
 Previous reviews focus exclusively on either a specific model architecture or an analysis task.
-For example, [8, 9, 10, 11] reviews deep learning methods for specific time series analysis tasks while failing to include advanced architecture such as Transformer.
-Several surveys [12, 13] provide up-to-date reviews for time series analysis focusing on specific deep learning architectures(i.e., Graph Neural Network and Transformer).
+For example, [8, 9, 10, 11] each review deep learning methods for specific time series analysis tasks while failing to include advanced architecture such as Transformers.
+Several surveys [12, 13] provide up-to-date reviews for time series analysis focusing on specific deep learning architectures (e.g., Graph Neural Networks and Transformers).
 Recently, BasicTS [14] and TFB [15] introduce forecasting benchmarks that enable an unbiased evaluation of existing approaches but do not provide an overview of the architectural design of those deep models.
 In this survey, we provide a comprehensive review of deep time series models for researchers and practitioners, starting from the basic modules to modern architectures.
-To foster practical applications, a time series benchmark is offered for a fair evaluation and identifying the effective scope of existing models.
-Our survey is organized as follows.
+To foster practical applications, a time series benchmark is offered for a fair evaluation and to identify the effective scope of existing models.
+This survey is organized as follows.
 Section 2 provides the background concepts of time series analysis.
 Section 3 introduces the basic modules that are widely utilized in prevalent deep time series models.
-Section 4 reviews the existing deep time series models in terms of the architecture design.
-Section 5 introduces the proposed open-source benchmark—Time Series Library (TSLib)—and presents extensive experimental comparison with detailed analysis.
+Section 4 reviews existing deep time series models in terms of the architecture design.
+Section 5 introduces an easy-to-use open-source benchmark, Time Series Library (TSLib), along with extensive experimental comparisons and detailed analyses.
 Section 6 provides a brief discussion of future research directions while Section 7 summarizes this survey.
@@
 2.1 Time Series
-Time series is a sequence of observations ordered by time, which can be denoted as , where represents the observed values at time point and is the number of variables.
-Since time series data are physical measurements obtained from sensors, systems are often recorded with multiple variables.
-Consequently, real-world time series usually recorded in a multivariate form.
-Theoretical studies [25, 26] have shown that when there are two or more non-stationary series, a linear combination of them can be stationary.
-This co-integration property helps in uncovering and modeling long-term relationships among non-stationary series.
-Therefore, the essence of time series analysis is to capture and utilize the temporal dependencies and inter-variable correlations within the observations.
-Temporal Dependency Given the sequential nature inherent in the observations, one evident technological paradigm is to capture the temporal dependence of a set of historical data.
-The basic idea of temporal dependencies is the intricate correlations between time points or sub-series.
-Traditional statistical models have laid the groundwork for modeling temporal dependencies.
-Prominent models include ARIMA (Autoregressive Integrated Moving Average) [1] have been extensively studied for capturing complex temporal patterns in the time series modality.
-Owing to their simplicity and interpretability, these statistical methods remain popular for tasks where the underlying temporal dynamics do not exhibit high complexity.
-Considering the high-dimensionality and non-stationarity of real-world time series, the research focus shifted towards deep learning for time series analysis.
-These advanced methods are designed to handle more complex temporal dynamics and offer greater flexibility in capturing the temporal dependency of time series data.
-Variate Correlation In addition to capturing temporal dependencies, understanding the variate correlations within high-dimensionality plays a pivotal role in analyzings multivariate time series.
-These correlations refer to the complex interactions and associations among different variables changing across the time.
-They provide valuable insights into the underlying dynamics and dependencies among the measurements, enabling a more comprehensive understanding of the latent process.
-Traditional approaches, such as Vector Autoregressive (VAR) models [27], extend the concept of autoregression to multiple variables and can capture the relationships between multiple quantities as they evolve over time.
-Technically, VAR represents each variable as a linear combination of its lagged values and the lagged values of all other variables in the model, which results in an inability to capture complex and non-linear relationships.
-Recently, advanced deep models, such as Graph Neural Networks [28] and Transformers [29, 30], have also been introduced for variate correlation modeling.
+Time series is a sequence of observations ordered by time, which can be denoted as , where represents the observed values at timestamp and is the number of variables.
+Note that spatiotemporal data (e.g.
+videos or weather records) can also be viewed as a special type of time series, where the “variate dimension” of the above notion corresponds to spatial pixels [25] in a video.
+However, previous research on spatiotemporal data primarily relies on highly structured spatial information [26], which limits their applications to general variates, like heterogeneous variates with distinct semantic meanings.
+In this paper, we focus on the more generic formalization, where we do not assume any spatial structure inside the variate dimension.
+Under this definition, the essence of time series analysis is to capture and utilize the temporal dependencies and inter-variable correlations within the observations.
+Temporal Dependency Given the inherent sequential nature of time series data, one evident technological paradigm is to capture the temporal dependence of a set of historical data.
+Essentially, temporal dependencies refer to the intricate connections that manifest among distinct time points or sub-series.
+Traditional statistical methods, such as those based on autoregressive principles represented by ARIMA [1], have been widely studied to model these dependencies.
+Owing to their simplicity and interpretability, these statistical methods retain prominence in scenarios where the underlying temporal dynamics exhibit limited complexity.
+Variate Correlation Beyond temporal dependencies, comprehending the correlations among distinct variables is paramount for analyzing multivariate time series, particularly in high-dimensional contexts.
+These correlations encapsulate the complex interactions and associations among multiple measurements as they evolve over time.
+Elucidating these relationships provides valuable insights into the underlying dynamics and latent processes.
+Conventional approaches, exemplified by VAR [27], extend the concept of autoregression to multiple variables, thereby enabling the capture of inter-variable relationships across time.
+However, VAR models represent each variable as a linear combination of lagged values of all other variables, which inherently limits their ability to capture misaligned or nonlinear relationships.
+The aforementioned fundamental elements of time series present considerable challenges for constructing practical analysis models.
+While traditional statistical methods offer foundational theoretical insights, their inherent reliance on presumptions of the data, such as linearity and stationarity, often restricts their applicability in intricate real-world scenarios.
+Recently, deep learning models have emerged as powerful alternatives for time series modeling.
+Their capacity for representation learning and scalability to big data uniquely positions them for high-dimensional and nonlinear time series data, effectively addressing diverse analysis tasks.
 2.2 Time Series Analysis Tasks
-Based on the understanding of underlying patterns and trends within time series data, time series analysis encompasses various downstream applications, including forecasting [31, 32], imputation [33, 34, 35], classification [8, 36], and anomaly detection [37, 5], each serving distinct purposes in diverse application domains.
-We illustrate representative time series analysis tasks in Figure 1.
-Forecasting is a fundamental task in time series analysis that requires models to uncover temporal dependencies and dynamic patterns in the data.
+Based on the understanding of underlying patterns and trends within time series data, time series analysis encompasses various downstream applications, including forecasting [28, 29], imputation [30, 31, 32], classification [8, 33], and anomaly detection [34, 5], each finding distinct applications in diverse domains, as illustrated in Figure 1.
+Forecasting is a fundamental task in time series analysis that requires models to uncover temporal dependencies and dynamic patterns.
 By capturing the relationships between past and future data, the forecasting model aims to predict future values or trends of the input series.
-Missing data due to sensor failures, data corruption, or absent measurements are ubiquitous in practical applications, leading to a growing demand for time series imputation to obtain higher-quality data.
+Missing data due to sensor failures, data corruption, or absent measurements is ubiquitous in practical applications, leading to a growing demand for time series imputation to obtain higher-quality data.
 Unlike forecasting, which predicts future values based on historical observations, imputation focuses on reconstructing missing values using the available contextual information.
@@
 Lastly, classification assigns a label or category to a given time series based on its characteristics, a task widely utilized in fields such as medical diagnosis.
+Given the pluralism of time series analysis, an array of deep learning models have emerged to address specific challenges in different tasks.
+Instead of a task-centric view into exhaustive details, this survey takes a model-centric view of deep time series models, focusing on their basic modules and architectural designs for addressing the foundational problems of capturing both temporal and inter-variable dependencies underlying large-scale time series datasets.
 3 Basic Modules
-Time series modeling approaches have evolved significantly, transitioning from traditional statistical models to sophisticated deep learning models.
+In the past decade, time series modeling approaches have evolved significantly, shifting from traditional statistical models to sophisticated deep learning models.
 Despite these advancements, many classical tools and analytical algorithms remain widely used and continue to serve as foundational design principles in modern deep models.
@@
 3.1 Stationarization
-As a foundational concept in time series analysis, stationarity refers to the property of a time series where its statistical properties remain constant over time.
-A stationary time series has a constant mean and variance, which simplifies statistical analysis and makes it easier to capture the underlying patterns and behavior within a time series.
-Since many statistics-based time series analysis methods take stationarity as a basic assumption, stationarization of time series data has become an essential module.
-There are ways of transforming non-stationary time series into stationary.
-Traditional time series models stationarize the time series through differencing or log-transformation.
+Stationarity is a foundational concept in time series analysis, referring to the property where the statistical properties of a time series remain constant over time.
+A stationary time series exhibits a constant mean and variance, which simplifies statistical analysis and makes it easier to capture the underlying patterns and behaviors within a time series.
+Given that many statistics-based time series analysis methods fundamentally presuppose stationarity, the transformation of non-stationary data into a stationary form has become an essential preprocessing step.
+Conventionally, this stationarization is achieved through methods such as differencing or log-transformation.
 In recent deep learning approaches, data normalization [35] takes the role of stationarization in a simple but effective way, which standardizes the value distribution of observations while maintaining the intrinsic variations and further helps mitigate the distribution shift between the source and target domains.
-The deep adaptive input normalization (DAIN) layer [39] was proposed to adaptively stationarize time series data according to their original distribution.
-RevIN [40] introduces reversible instance normalization to time series data, which is an effective normalization-and-denormalization method with learnable affine transforms to make the model bypass the non-stationary inputs.
-Non-Stationary Transformer [41] (Stationary for short in the following) proposes a simpler but more effective series stationarization technique that improves the predictive capability of non-stationary series without extra parameters.
-Specifically, for a sequence with time stamps and variates , the outline of Stationary [41] can be summarized as:
-where is in a small value for numerical stability.
+Recent deep learning architectures have developed specialized layers to address the non-stationary problem.
+The deep adaptive input normalization (DAIN) layer [36] was proposed to adaptively standardize time series data based on its original distribution.
+RevIN [37] introduces a reversible instance normalization technique with learnable affine transforms to make the model bypass the non-stationary inputs.
+Further, Non-Stationary Transformer [38] (Stationary for short in the following) proposes a simpler but more effective series stationarization technique that improves the predictive capability of non-stationary series without extra parameters.
+Specifically, for a sequence with time stamps and variates , where , the outline of Stationary [38] can be summarized as follows:
+where is a small value for numerical stability.
 are the variate-specific mean and variance.
-To recover the distribution and non-stationarity of the original series, a de-normalization module is further used to augment the model output with mean and variance statistics of inputs.
-The idea of stationarization and the above-mentioned techniques have been widely used in subsequent deep time series models [23, 30, 42, 43].
-Recent SAN [44] rethinks the nature of non-stationary data and tries to split it into non-overlap equally-sized slices and perform normalization on each slice.
-Specifically, based on the evolving trends of statistical properties, SAN introduces a statistics prediction module to predict the distributions of future slices.
+To recover the distribution and non-stationarity of the original series, a denormalization module is further used to augment the model output with mean and variance statistics of inputs.
+The idea of stationarization and the aforementioned techniques have seen widespread adoption in subsequent deep time series models [23, 39, 40, 41].
+For instance, recent works such as SAN [42] rethink the nature of non-stationary data by normalizing non-overlapping, equally-sized slices instead of the entire series, and incorporating a statistics prediction module to forecast future slice distributions.
 3.2 Decomposition
-Decomposition [45, 46], as a conventional approach in time series analysis, can disentangle time series into several components with categorized patterns, and works primarily useful for exploring complex series variations.
-In the previous work, diverse decomposition paradigms are explored.
+Decomposition [43, 44], as a conventional approach in time series analysis, breaks time series into several components of more regular patterns, with a divide-and-conquer paradigm to explore complex series variations.
+In the previous work, a set of typical decomposition schemes are explored.
 3.2.1 Seasonal-Trend Decomposition
-Seasonal-trend decomposition [47, 48] is one of the most common practices to make raw data more predictable, which can separate the series into several different components: trend, seasonal, cyclical, and irregular, namely
-where the trend component represents the overall long-term pattern of the data over time, the cyclical component reflects repeated but non-periodic fluctuations within data, the seasonal component indicates the repetitive patterns over a fixed period, and the irregular component is the residuals or remainder of the time series after the other components have been removed.
-The trend-seasonality decomposition can be achieved by using mathematical tools such as filters or exponential smoothing [49, 50].
+Seasonal-trend decomposition [45, 46] is one of the most common practices to enhance the predictability of raw time series data.
+This technique separates a series into distinct additive components: trend, seasonal, cyclical, and irregular.
+The trend component represents the overall long-term pattern of the data over time, the cyclical component reflects repeated but non-periodic fluctuations within the data, the seasonal component indicates the repetitive patterns over a fixed period, and the irregular component is the residual or remainder of the time series after the other components have been removed.
+Such decomposition can be achieved using mathematical tools like filters or exponential smoothing [47, 48].
 Previous statistical approaches mainly adopt the trend-seasonality decomposition as data pre-processing [49].
-In deep models, Autoformer [22] firstly introduces the idea of decomposition to deep learning architecture and proposes a series decomposition block as a basic module to extract the seasonal and trend-cyclical parts of deep features and input series, whose computation process can be formalized as:
-The series decomposition block is concisely implemented based on a temporally average pooling layer with padding operation to keep the sequence length unchanged.
-This design can capture trends , and the remainder is taken as the seasonal part .
-The proposed series decomposition block has been widely used in the follow-up [6, 52, 53, 54, 55] as a native building block of deep models to disentangle the underlying patterns of deep features.
+In deep models, Autoformer [22] firstly introduces the idea of decomposition to deep learning architecture and proposes a series decomposition block as a basic module based on the average pooling to extract seasonal and trend-cyclical parts of deep features and input series:
+Implemented based on a temporally average pooling layer, the proposed decomposition block effectively captures the trends , and the remainder is taken as the seasonal part .
+This design has been widely adopted as a native building block in the subsequent works [6, 50, 51, 52, 53] for disentangling underlying patterns within deep features.
 3.2.2 Basis Expansion
 Basis expansion is a mathematical method used to represent a function or a set of data points in terms of a new set of pre-defined functions.
-These new functions form a basis for a function space, meaning any function in that space can be expressed as a linear combination of these basis functions.
-In the context of time series analysis, basis expansion is used to reveal complex non-linear temporal relationships by decomposing the time series into a combination of basic variations, which also enhances interpretability.
+These new functions form the bases for a function space, allowing any function in that space to be expressed as a linear combination of these basis functions.
+In the context of time series analysis, basis expansion is used to reveal nonlinear temporal relationships by decomposing the time series into a combination of basic variations, thereby enhancing interpretability.
 As a representative model, N-BEATS [54] presents hierarchical decomposition to time series by utilizing a fully connected layer to produce expansion coefficients for both backward and forward forecasts.
-For -th blocks in the proposed hierarchical architecture, the operation can be as follows:
+For the -th block in the proposed hierarchical architecture, the operation is
 where is the backcast results which restrict the block to approximate the input signal , then removes the portion of well-estimated signal from , therefore providing a hierarchical decomposition.
 is the partial forecast based on the decomposed input and the final forecast is the sum of all partial forecasts.
-Subsequently, N-HiTs [57] redefine the N-BEATS by incorporating subsampling layers before the fully connected blocks, which enhances the input decomposition via multi-frequency data sampling and future predictor via multi-scale interpolation.
-DEPTS [58] puts forward a novel decoupled formulation for periodic time series by introducing the periodic state as a hidden variable and then develops a deep expansion module on top of residual learning to conduct layer-by-layer expansions between observed signals and hidden periodic states.
-Similarly, DEWP [59] is a also stack-by-stack expansion model to handle multivariate time series data, where each stack consists of a variable expansion block to capture dependencies among multiple variables and a time expansion block to learn temporal dependencies.
+Subsequently, N-HiTs [55] redefine the N-BEATS by incorporating subsampling layers before the fully connected blocks.
+This modification enhances input decomposition through multi-frequency data sampling and improves future prediction via multi-scale interpolation.
+DEPTS [56] proposes a novel decoupled formulation for periodic time series by introducing a periodic state as a hidden variable.
+It then develops a deep expansion module atop residual learning to perform layer-by-layer expansions between observed signals and these hidden periodic states.
+Similarly, DEWP [57] is a stack-by-stack expansion model designed to handle multivariate time series data.
+Each stack within the model comprises a variable expansion block for capturing dependencies among multiple variables and a time expansion block for learning temporal dependencies.
 3.2.3 Matrix Factorization
-The above-mentioned two decomposition methods are proposed for univariate series or applied to multivariate series in a variate-independent way.
-Here, we discuss a factorization-based decomposition for multivariate series.
-Specifically, many multivariate time series data in real-world scenarios can also be referred to as high-dimensional data.
-They can be formalized in the form of a matrix, whose rows correspond to variate and columns correspond to time points.
-Since variables in multivariate time series tend to be highly correlated, it can be possibly reduced to a more compact space.
-Matrix factorization methods [60] work by decomposing the high-dimensional series data into the product of two matrices in a lower-dimensional latent space.
-For a multivariate time series , as shown in Figure 2 the matrix can be approximated by the multiplications of two lower rank embedding matrix, , in which , and is a hyperparameter.
-Besides the estimation, there are regularizers to avoid overfitting problems in factorization.
+The above-mentioned two decomposition methods primarily address univariate series or are applied to multivariate series in a variate-independent way.
+Here, we explore factorization-based decomposition that is particularly pertinent to multivariate time series.
+These high-dimensional data can be formalized as a matrix where rows represent variates and columns denote time points.
+Since variables in multivariate time series tend to be highly correlated, they can be reduced to a more compact space.
+Matrix factorization methods [58] achieve this by decomposing the high-dimensional series data into the product of two matrices in a lower-dimensional latent space.
+For a multivariate time series , the matrix can be approximated by the multiplications of two lower rank embedding matrix, , in which , and is a hyperparameter.
+Besides the estimation, regularizers are critical in factorization-based models to mitigate overfitting.
 Going beyond the canonical design that takes the squared Frobenius norm as regularizers, temporal regularized matrix factorization (TRMF) [59] designs an autoregressive-based temporal regularizer to describe temporal dependencies among latent temporal embeddings.
 Further, [60] extended TRMF with a new spatial autoregressive regularizer to estimate low-rank latent factors by simultaneously learning the spatial and temporal autocorrelations.
-NoTMF [63] integrates the vector autoregressive process with differencing operations into the classical low-rank matrix factorization framework to better model real-world time series data with trend and seasonality.
-Eliminating the need for tuning regularization parameters, BTF [64] is a fully Bayesian model that integrates the probabilistic matrix factorization and vector autoregressive process into a single probabilistic graphical model.
-Instead of using an autoregressive-based temporal regularization, DeepGLO [65] utilizes a temporal convolution network for regularization to capture non-linear dependencies.
-LSTM-GL-ReMF [66] contains an LSTM-based temporal regularizer to learn complex long-term and short-term non-linear temporal correlations and a Graph Laplacian spatial regularizer [67] to capture spatial correlations.
+For enhanced modeling of the trends and seasonality of real-world time series data, NoTMF [61] integrates the vector autoregressive process with differencing operations into the classical low-rank matrix factorization framework.
+Addressing the challenge of parameter tuning, BTF [62] offers a fully Bayesian approach, uniting probabilistic matrix factorization and the vector autoregressive process within a single probabilistic graphical model.
+Diverging from autoregressive-based temporal regularization, DeepGLO [63] leverages a temporal convolution network to capture non-linear dependencies.
+Similarly, LSTM-GL-ReMF [64] employs an LSTM-based temporal regularizer for complex long-term and short-term non-linear temporal correlations, complemented by a graph Laplacian spatial regularizer [65] for capturing spatial correlations.
 3.3 Fourier Analysis
-Fourier analysis [68, 69] can convert a physical signal into the Fourier domain to highlight the inherent frequency properties of the original data and has been a well-acknowledged analysis tool in extensive areas.
-Since time series are usually recorded as a sequence of discrete time points by sampling the original continuous signals, Fourier analysis has become one of the mainstream tools in time series modeling and has been demonstrated favorable effectiveness and efficiency [70, 71].
-Introducing the Fourier domain not only augments the representation of the original series but also provides a global view since the frequency spectrum distribution, which can indicate essential periodic properties of time series.
-In practice, Fast Fourier Transform (FFT) [72] and Wavelet Transform (WT) [73] as the basic algorithms connecting the discrete temporal domain to the frequency domain, have gained increasing popularity in the modular design of deep time series models [74, 75, 76, 77, 78, 79, 80, 81].
-Existing approaches can be roughly divided into two categories: time-domain and frequency-domain modeling.
+Fourier analysis [66, 67] is a powerful mathematical technique that transforms a physical signal into the frequency domain, thereby illuminating the inherent periodic properties of the original data.
+It has long been recognized as a fundamental analytical tool across diverse domains.
+Given that time series data are typically recorded as sequences of discrete observations sampled from continuous signals, Fourier analysis has naturally emerged as a mainstream approach in time series modeling with favorable effectiveness and efficiency [68, 69].
+Introducing the Fourier domain not only enhances the representation of the original series but also provides a global perspective on temporal dynamics.
+The resulting frequency spectrum distribution offers crucial insights into the essential periodic components and underlying cyclical behaviors of time series.
+In practice, algorithms such as Fast Fourier Transform (FFT) [70] and Wavelet Transform (WT) [71] act as foundational bridges, connecting the discrete temporal domain to the frequency domain.
+These techniques have gained increasing popularity in the modular design of deep time series models [72, 73, 74, 75, 76, 77, 78, 79].
+Existing approaches can be broadly categorized into two main types: time-domain and frequency-domain modeling.
 3.3.1 Time-Domain Modeling
@@
 Consequently, it can be used to identify potentially dominant periods and their corresponding frequencies in the data by analyzing the highest amplitude components.
-As a typical practice, TimesNet [3] employs the Fast Fourier Transform (FFT) to extract the most significant frequencies with the highest amplitude values, subsequently reshaping the 1D time series data into a 2D space based on the identified periods for better representation learning.
-Following TimesNet, PDF [82] posits that frequencies with larger values facilitate a more discernible distinction between long-term and short-term relationships.
-In addition to exploiting the information of the sequence obtained by the Fourier Transformer, some works attempt to perform efficient computation through the Fast Fourier Transformer.
+As a prominent integration, TimesNet [3] employs the Fast Fourier Transform (FFT) to extract the most significant frequencies with the highest amplitude values.
+It then reshapes the 1D time series data into a 2D space based on these identified periods, thereby facilitating hierarchical representation learning.
+Building upon TimesNet, PDF [80] posits that frequencies with larger values facilitate a more discernible distinction between long-term and short-term relationships.
+The frequency-domain representation provides valuable information regarding both amplitudes and phases.
+Low-frequency components, for instance, correspond to slower variations or underlying trends in the signal, while high-frequency components capture finer details or rapid fluctuations.
+A significant body of work has focused on leveraging frequency-domain information to enhance the model’s capability in capturing temporal dependencies.
+FiLM [81] combines Fourier analysis with low-rank approximation and introduces a frequency enhanced layer (FEL).
+This approach retains portions of the representation related to low-frequency Fourier components and the top eigenspace, thereby effectively reducing noise and accelerating training speed.
+FITS [82] integrates a low-pass filter (LPF) to eliminate high-frequency components above a specified cutoff frequency, thereby compressing the model size while preserving essential information.
+Taking an opposing view, FEDformer [83] posits that retaining only low-frequency components is insufficient for time series modeling, as it risks dismissing important data fluctuations.
+Based on this consideration, FEDformer aims to capture a global view of time series by representing the series using a randomly selected constant number of Fourier components, encompassing both high-frequency and low-frequency contributions.
+Beyond merely exploiting sequence information obtained from Fourier transforms, some research focuses on achieving efficient computation through the Fast Fourier Transform.
 Auto-correlation is a fundamental concept in time series analysis that measures the dependence between observations at different time points within a sequence of data.
 The Wiener-Khinchin theorem [84] provides a mathematical relationship between the auto-correlation function and the power spectral density (PSD) of a stationary random process, where the auto-correlation function represents the inverse Fourier transform of the PSD.
-Taking the data as a real discrete-time process, Autoformer [22] proposes an Auto-Correlation mechanism with an efficient Fast Fourier Transforms to capture the series-wise correlation.
-The frequency-domain representation provides information about the amplitudes and phases, where low-frequency components correspond to slower variations or trends in the signal, and high-frequency components capture fine details or rapid variations.
-A significant body of work has focused on leveraging frequency-domain information to enhance the model’s capability in capturing temporal dependencies.
-FiLM [84] introduces Frequency Enhanced Layers (FEL) which combine Fourier analysis with low-rank approximation to keep the part of the representation related to low-frequency Fourier components and the top eigenspace to effectively reduce the noise and boost the training speed.
-FITS [85] integrates a low-pass filter (LPF) to eliminate high-frequency components above a specified cutoff frequency, thereby compressing the model size while preserving essential information.
-From an opposite idea, FEDformer [86] posits that retaining only low-frequency components is insufficient for time series modeling, as it may dismiss important fluctuations in the data.
-Based on the above considerations, to capture the global view of time series, FEDformer represents the series by randomly selecting a constant number of Fourier components, including both high-frequency and low-frequency components.
+Taking the data as a real discrete-time process, Autoformer [22] proposes an Auto-Correlation mechanism with an efficient Fast Fourier Transform to capture the series-wise correlation.
 3.3.2 Frequency-Domain Modeling
-Building on time-frequency analysis in signal processing, several approaches have been developed to study time series simultaneously in both the time and frequency domains.
-ATFN [87] comprises an augmented sequence-to-sequence model that learns the trending features of complex non-stationary time series, along with a frequency-domain block designed to capture dynamic and intricate periodic patterns.
-TFAD [88] introduces a time-frequency analysis-based model that employs temporal convolutional networks to learn both time-domain and frequency-domain representations.
+Building on advancements in time-frequency analysis from signal processing, several approaches have been developed to analyze time series simultaneously in both the time and frequency domains.
+ATFN [85], for instance, utilizes an augmented sequence-to-sequence model to learn trending features of non-stationary time series, alongside a frequency-domain block for capturing periodic patterns.
+TFAD [86] introduces a time-frequency analysis-based model that employs temporal convolutional networks to obtain both time-domain and frequency-domain representations.
 Some works have developed specialized deep learning architecture to process the frequency domain of time series.
@@
 FCVAE [90] integrates both the global and local frequency features into the condition of Conditional Variational Autoencoder (CVAE) concurrently.
-Recent TSLANet [92] propose a lightweight Adaptive Spectral Block (ASB) to replace the self-attention mechanism, which is achieved via Fourier-based multiplications by global and local filters.
-FourierDiffusion [93] explores extending the score-based SDE formulation of diffusion to complex-valued data and therefore implements time series diffusion in the frequency domain.
+Recent TSLANet [91] proposes a lightweight Adaptive Spectral Block (ASB) to replace the self-attention mechanism, which is achieved via Fourier-based multiplications by global and local filters.
+FourierDiffusion [92] extends the score-based SDE formulation of diffusion models to complex-valued data and therefore implements time series diffusion in the frequency domain.
 4 Model Architectures
-As we have discussed in Section 2, the time series model needs to unearth the intrinsic temporal dependencies and variate correlations that lie in observations.
-In this section, we provide a technical review of the existing deep time series models.
-As we have presented in Figure 3, existing works can be classified into five categories based on their backbone architecture, namely MLP-based, RNN-based, CNN-based, GNN-based, and Transformer-based.
-4.1 Multi-Layer Perceptrons
-As a representation of traditional statistical time series models, the Auto-regressive (AR) model assumes that the model output depends linearly on its own historical values.
-Inspired by the remarkable performance of auto-regressive models, Multi-Layer Perceptrons (MLP) have become a popular architecture for modeling time series data.
-As a representative work of linear-based models, N-BEATS [24] is a pure MLP-based deep time series model without any time-series-specific knowledge to capture the temporal patterns in time series.
-Specifically, as described in Equ.
-(4) N-BEATS consists of deep stacks of fully-connected layers with two residual branches in each layer, one is for the backcast prediction and the other one is the forecast branch.
-Extending the idea of neural basis expansion analysis, N-HiTs [57] use multi-rate signal sampling and hierarchical interpolation and N-BEATSx [94] incorporate exogenous variables to enhance the prediction.
-Recent research by DLinear [52], also referred to as LTSF-Linear, challenges the effectiveness of complicated deep architecture in temporal modeling.
+As discussed in Section 2, time series models aim to unearth the intrinsic temporal dependencies and variate correlations to uncover underlying patterns and dynamics within observations.
+Existing approaches can be broadly categorized into five groups based on their architectural backbones (Figure 2), namely, MLP-based, RNN-based, CNN-based, GNN-based, and Transformer-based.
+Each of these architectures possesses distinct strengths and design principles that address specific challenges in time series analysis.
+In this section, we provide a comprehensive technical review of current deep time series models, specifically highlighting their distinct approaches to time series modeling.
+4.1 Multilayer Perceptrons
+As a cornerstone of traditional statistical time series models, the Auto-regressive (AR) model posits that the model output depends linearly on its own past values.
+Inspired by the notable efficacy of auto-regressive models, Multilayer Perceptrons (MLP) have become a popular architecture for time series modeling.
+As a prime example of MLP-based models, N-BEATS [24] integrates no time-series-specific components to capture temporal patterns.
+Specifically, as described in Eq.
+(3), N-BEATS consists of deep stacks of fully-connected layers with two residual branches in each layer, one for the backcast prediction and the other for the forecast branch.
+Expanding the idea of neural basis expansion, N-HiTs [55] enhances prediction through multi-rate signal sampling and hierarchical interpolation, while N-BEATSx [93] further extends the framework by incorporating exogenous variables.
+Recent research, notably DLinear [50], rethinks the effectiveness of complicated deep architectures for temporal modeling.
 It argues a simple linear regression in the raw space that achieves remarkable performance in both modeling and efficiency.
-As illustrated in Figure 4, prevalent MLP-based deep time series models consist of simple linear layers primarily designed for forecasting tasks.
-Also lightweight but effective, FITS [85] advocates time series analysis can be treated as interpolation exercises within the complex frequency domain and further introduces a complex-valued linear layer to learn amplitude scaling and phase shift in the frequency domain.
-Inspired by MLP-Mixer [95] in computer vision, several works have attempted to utilize MLPs to model both temporal and variate dependencies.
+As illustrated in Figure 3, many prevalent MLP-based deep time series models, primarily designed for forecasting, rely on such simple linear layers.
+Also lightweight but effective, FITS [82] advocates that time series forecasting can be treated as interpolation exercises within the complex frequency domain.
+It introduces a complex-valued linear layer to learn amplitude scaling and phase shift directly in the frequency domain.
+Inspired by MLP-Mixer [94] in computer vision, several works have attempted to leverage MLPs to model both temporal and variate dependencies.
 TSMixer [95] contains interleaving time-mixing and feature-mixing MLPs to extract information from different perspectives.
-To better model the global dependencies in time series data, FreTS [90] investigates the learned patterns of frequency-domain MLPs which are operated on both inter-series and intra-series scales to capture channel-wise and time-wise dependencies in multivariate data.
-Recent works have moved beyond using simple linear layers over discrete time points.
-TimeMixer suggests that time series exhibit distinct patterns in different sampling scales and proposes an MLP-based multiscale mixing architecture.
-TiDE [97] incorporates exogenous variables to enhance the time series prediction.
+To better capture global dependencies, FreTS [89] investigates the learned patterns of frequency-domain MLPs, which operate on both inter-series and intra-series scales to capture channel-wise and time-wise dependencies in multivariate data.
+Moving beyond simple linear layers over discrete time points, TimeMixer suggests that time series exhibit distinct patterns across different sampling scales and introduces an MLP-based multiscale mixing architecture.
+TiDE [96] incorporates exogenous variables to enhance time series prediction.
 Based on Koopman theory and Dynamic Mode Decomposition (DMD) [97], which is a dominant approach for analyzing complicated dynamical systems, Koopa [98] hierarchically disentangles dynamics through an end-to-end predictive training framework and can utilize real-time incoming series for online development.
 4.2 Recurrent Neural Networks
-Recurrent Neural Networks (RNNs) are specifically designed to model sequential data [100, 101, 102], such as natural language processing [103] and audio modeling [104].
-Since time series are also serial in nature, RNNs have emerged as a popular choice for analyzing time series data [105].
-Existing RNN-based deep time series models focus on combating the gradient vanishing problem caused by the vanilla recurrent structure and modeling the mutual correlation among multivariate variables.
-Previous works [106, 107, 108, 109, 110] use variants of RNN to model temporal dependencies.
+Recurrent Neural Networks (RNNs) are specifically designed for sequential data [99, 100, 101], such as natural language processing [102] and audio modeling [103].
+Given the inherent sequential nature of time series, RNNs have naturally become a popular choice for time series analysis [104].
+Existing RNN-based deep time series models primarily focus on combating the gradient vanishing problem caused by the vanilla recurrent structure and modeling the mutual correlation among multivariate variables.
+Numerous prior works [105, 106, 107, 108, 109] have leveraged various RNN variants to capture complex dependencies.
 LSTNet [110] combines the recurrent structure with the convolutional layer to capture both the short-term local dependency between variables and long-term patterns for time series.
-Moreover, a novel recurrent-skip component based on the periodic pattern is introduced to alleviate gradient vanishing in modeling long-term dependencies.
+It further introduces a novel recurrent-skip component, based on periodic patterns, to alleviate gradient vanishing in modeling long-term dependencies.
 Similarly, DA-RNN [111] combines the recurrent unit with a dual-stage attention mechanism to adaptively extract relevant series at each time step.
@@
 Technologically, it learns not only the seasonal behavior with time series but dependencies on given covariates across time series, allowing the model to make predictions even when there is little or no historical data.
-Also based on Markovian state representation, the State Space Model (SSM) [114] is another classical mathematical framework that captures the probabilistic dependence between observed measurements in stochastic dynamical systems.
-Concretely, a single-input single-output (SISO) linear state space model is defined as follows:
-where are input signal, latent state, and output signal respectively. The system is characterized by the matrices can be learned by the deep neural network.
-SSMs have proven their effectiveness and efficiency in processing well-structured time series data, but traditional approaches have to refit each time series sample separately and therefore cannot infer shared patterns from a dataset of similar time series.
-With the rise of deep learning models, modern SSMs are often implemented in a recurrent manner.
-By adapting and propagating a deterministic hidden state, RNNs are able to represent long-term dependencies in continuous data which offer an alternative to classical state space models.
-Therefore, some work [115, 116] have attempted to fuse classical state space models with deep neural networks.
-Representative like Deep State Spaces Model (DSSM) [117], using a recurrent neural network (RNN) to parametrize a particular linear SSM, takes advantage of incorporating structural assumptions and learning complex patterns.
-Structured State Space sequence model (S4) [118] introduces a new parameterization for the SSM by conditioning matrix with a low-rank correction, allowing it to be diagonalized stably, which empowers the model with better long-term modeling capacity.
-Similar to S4, LS4 [119] is a generative model with latent space evolution following a state space ordinary differential equations (ODE).
-Recent work on Mamba [120] has emerged as a powerful method for modeling long-context sequential data while scaling linearly with sequence length.
-Utilizing a simple selection mechanism that parameterizes the SSM parameters based on the input, Mamba can discern the importance of information in a manner similar to the attention mechanism, posing a potentially effective way to sequential modeling.
+Recent work [77] introduces a novel DTWAtt module, which integrates a data-dependent wavelet function into the BiLSTM network as wavelet attention to analyze dynamic frequency components in nonstationary time series.
+Based on Markovian state representation, the State Space Model (SSM) [113] is a classical mathematical framework for capturing probabilistic dependencies between observed measurements in stochastic dynamical systems.
+While SSMs have proven their effectiveness and efficiency in processing well-structured time series data, traditional approaches have to refit each series separately and therefore cannot infer shared patterns from a dataset of similar time series.
+With the advent of deep learning, modern SSMs are often implemented in a recurrent manner.
+By adapting and propagating a deterministic hidden state, RNNs offer an alternative to classical state space models by representing long-term dependencies in continuous data.
+Consequently, some works [114, 115] have sought to fuse classical state space models with deep neural networks.
+Representative among these, the Deep State Space Model (DSSM) [116] leverages a recurrent neural network (RNN) to parametrize a specific linear SSM, thereby incorporating structural assumptions while learning complex patterns.
+The Structured State Space sequence model (S4) [117] introduces a novel parametrization for the SSM by conditioning matrix with a low-rank correction, which allows for stable diagonalization and significantly enhances the model’s long-term modeling capacity.
+More recently, Mamba [118] has emerged as a powerful method for modeling long sequential data with linear complexity w.r.t.
+sequence length.
+Utilizing a simple selection mechanism that parameterizes the SSM parameters based on input, Mamba can discern valuable information in observations, posing an effective way for time series analysis.
+Going beyond learning among discretized time steps, Neural Ordinary Differential Equations (Neural ODEs) [119] extend RNNs into a continuous-time version.
+Technically, Neural ODEs parameterize the derivative of hidden states through neural networks, enabling seamless modeling of temporal evolution in a differentiable manner.
+Given the ability to model continuous-time, several works have explored applying Neural ODEs to handle irregularly-sampled time series data [120, 121].
+Building on these advancements, [122] integrates Neural ODEs with physical priors to augment physical models with deep networks for forecasting complex dynamics.
+These capabilities make Neural ODEs a compelling complement to RNN-based methods, offering a more flexible and robust approach for time series analysis.
 4.3 Convolutional Neural Networks
-Since the semantic information of time series is mainly hidden in the temporal variation, Convolutional neural networks (CNN) [18, 121] have become a competitive backbone for their ability to capture local features and pattern recognition.
-By leveraging convolutions and hierarchical feature extraction, CNNs have shown remarkable success in various computer vision tasks, such as image classification [122], segmentation [123] and object detection [124].
-Considering the temporal continuity of time series data, previous works [125, 126, 127] apply one-dimensional CNN (1D CNN) to capture the local patterns of time series data.
-Recent SCINet[128] applies normal convolutions with a hierarchical downsample-convolve-interact architecture to capture dynamic temporal dependencies at different temporal resolutions of time series data.
+Since the semantic information of time series is mainly hidden in the temporal variation, Convolutional neural networks (CNN) [18, 123] have become a competitive backbone.
+Their strength lies in their ability to capture local features and recognize intricate patterns, which has been demonstrated with remarkable success in various computer vision tasks, including image classification [124], segmentation [125], and object detection [126].
+Considering the temporal continuity of time series data, previous works [127, 128, 129] apply a one-dimensional CNN (1D CNN) to discern the local patterns of time series data.
+Recent SCINet[130] employs canonical convolutions with a hierarchical downsample-convolve-interact architecture to capture dynamic temporal dependencies across varying temporal resolutions.
 Inspired by the idea of masked convolution [25], Wavenet[131] introduces causal convolution and dilated causal convolution to model long-range temporal causality.
-Similar to Wavenet, Temporal Convolutional Networks (TCN) [131] uses a stack of dilated convolutional kernels with progressively enlarged dilation factors to achieve a large receptive field.
-However, the limited receptive field of TCN makes it difficult for them to capture global relationships in time series data.
+Similarly, Temporal Convolutional Networks (TCN) [132] uses a stack of dilated convolutional kernels with progressively enlarged dilation factors to achieve a large receptive field.
+However, the limited receptive field of TCN often hinders its capacity to capture global relationships in time series data.
 Based on TCN, MICN[133] is a local-global convolution network that combines different convolution kernels to model temporal correlation from a local and global perspective.
-ModernTCN [133] boosts the traditional TCN to capture cross-time and cross-variable dependency by DWConv and ConvFFN separately.
-Considering that DWConv is proposed to learn temporal information, it is operated variate-independently to learn the temporal dependency of each univariate time series.
-Beyond 1D space, motivated by the periodicity properties of time series data, TimesNet [3] transforms the 1D time series data into a set of 2D tensors in each TimesBlock based on the estimated period lengths, where the inter-period variations are presented in tensor columns and inner-period ones are shown in tensor rows.
+ModernTCN [134] enhances the traditional TCN by separately employing Depthwise Separable Convolutions (DWConv) and Convolutional Feed-Forward Networks (ConvFFN) to capture cross-time and cross-variable dependencies, respectively.
+Motivated by the inherent periodicity of time series data, TimesNet [3] goes beyond traditional 1D modeling.
+It [3] transforms the 1D time series data into a set of 2D tensors in each TimesBlock based on the estimated period lengths, where the inter-period variations are presented in tensor columns and inner-period ones are shown in tensor rows.
 Here is a hyperparameter, corresponding to multiple 1D-to-2D transformations with different periods.
-Then it applies inception block [134, 135] to process the transformed 2D tensors, which can be summarized as:
+Subsequently, TimesNet applies an Inception Block [135, 136] to process these transformed 2D tensors, which can be summarized as:
 where is the -th transformed 2D tensor.
@@
 4.4 Graph Neural Networks
-Analyzing multivariate time series data is often challenging due to the complex and often non-linear correlations between variables.
-To address this challenge, Graph neural networks (GNNs) [136, 137] have been widely adopted in time series analysis.
-By modeling multivariate data as a spatiotemporal graph, where each node represents a variable, GNNs can extract relationships among neighboring nodes and capture the temporal evolution of node attributes over time, thereby providing a robust framework for understanding the underlying dynamics of multivariate time series.
-The core goal of GNN architecture is to model the underlying topological relations in multivariate data, therefore existing GNN-based works can be roughly divided into two categories based on whether graph structure is part of the input into the model.
+Analyzing multivariate time series data is often challenging due to the complicated and often nonlinear correlations between variables.
+To address this, Graph neural networks (GNNs) [137, 138] have been widely adopted in time series analysis.
+As depicted in Figure 6, by modeling multivariate data as a spatiotemporal graph, where each node represents a variable, GNNs can extract relationships among neighboring nodes and capture the temporal evolution of node attributes over time, providing a framework for understanding the underlying dynamics of multivariate time series.
+The primary goal of GNN architecture is to model the latent topological relations in multivariate data.
+Accordingly, existing GNN-based works can be broadly categorized into two types: those that rely on a pre-defined graph structure as input and those that learn the graph structure dynamically.
 DCRNN [139] models the spatial dependency of traffic as a diffusion process on a directed graph and uses diffusion convolution to capture the spatial dependency, alongside a recurrent neural network to capture temporal dynamics.
 Similarly, STGCN [140] integrates graph convolutional networks to model the spatial dependencies among traffic sensors with temporal convolutions to capture the temporal dependencies in the traffic time series data.
-Graph WaveNet [140] combines graph convolution with dilated casual convolution and learns an adaptive dependency matrix through node embedding, enabling the model to automatically capture hidden spatial dependencies in spatial-temporal graph data.
+Graph WaveNet [141] combines graph convolution with dilated causal convolution and learns an adaptive dependency matrix through node embedding, enabling the model to automatically capture hidden spatial dependencies in spatial-temporal graph data.
 Similarly, AGCRN [142] enhances the traditional graph convolutional network with node adaptive parameter learning and data-adaptive graph generation modules, allowing for the automatic capture of spatial and temporal correlations without a pre-defined graph structure.
@@
 4.5 Transformers
-In the view of the great success in the field of natural language processing [16, 144, 145, 146, 147] and computer vision [19, 148, 149, 150], Transformers have also emerged as a powerful backbone for time series analysis.
-Benefiting from the self-attention mechanism [151], Transformer-based models can capture long-term temporal dependencies and complex multivariate correlations.
+In view of the great success in the field of natural language processing [16, 145, 146, 147, 148] and computer vision [19, 149, 150, 151], Transformers have emerged as a powerful backbone for time series analysis.
+Benefiting from the self-attention mechanism [152], Transformer-based models excel at capturing long-term temporal dependencies and entangled multivariate correlations.
 As overviewed in Figure 7, existing Transformer-based time series models can be categorized based on the granularity of representation used in the attention mechanism, namely point-wise, patch-wise, and series-wise approaches.
@@
 4.5.1 Point-wise Dependency
-Due to the serial nature of time series, most existing Transformer-based works use a point-wise representation of time series data and apply attention mechanisms to capture the correlations among different time points.
-Among these point-wise modeling approaches, Data Embedding is a crucial component that maps the value of time series data to a high-dimensional representation.
-Given time series with corresponding time stamp information , where is the variate number and is the types of time stamps, the embedding module can be summarized as follow:
+Due to the serial nature of time series, most existing Transformer-based works learn point-wise representations and apply attention mechanisms to capture the correlations among different time points.
+Among these point-wise modeling approaches, Data Embedding stands as a crucial component, mapping raw time series values into a high-dimensional representation.
+Given time series with corresponding time stamp information , where is the variate number and is the types of time stamps, the embedding module can be summarized as:
 where and is the dimension of the emebedded representation, value projection and timestamp embedding are implemented by channel-dimension linear layers, and denotes the absolute position embedding to preserve the sequential context of input series.
-To better apply the Transformer architecture to the time series domain, researchers have explored two aspects: designing pre-processing modules and modifying the attention mechanisms.
-As we have discussed in Section 3.1, past RevIN [40] and Stationary [41] achieved superior performance by introducing Normalization and De-Normalization modules before and after the vanilla Transformer.
-Besides, Stationary [41] further proposes De-stationary Attention to avoid the over-stationarization problem.
-Given that the canonical attention approach leads to a quadratic computational complexity, numerous efficient Transformers [21, 22, 153, 86] have been proposed to mitigate the complexity caused by point-wise modeling, which is summarized in Table II.
+To optimize the application of Transformer architecture to the time series domain, existing literature primarily falls into two categories: designing pre-processing modules and modifying the attention mechanisms.
+As previously discussed in Section 3.1, techniques like RevIN [37] and Stationary [38] have achieved superior performance by integrating Normalization and De-Normalization modules before and after the Transformer blocks.
+Notably, Stationary [38] further introduces De-stationary Attention to mitigate the problem of over-stationarization.
+Given that canonical attention leads to a quadratic computational complexity, various efficient attentions [21, 22, 154, 83] have been proposed to mitigate the complexity arising from point-wise modeling, which is summarized in Table II.
 LogSparse [153] proposes Convolutional Self-Attention to replace canonical attention by employing causal convolutions to produce queries and keys in the self-attention layer.
-Informer [21] introduces a Query Sparsity Measurement, where a larger value indicates a higher chance to contain the dominant information in self-attention.
-Based on the proposed sparsity measurement, it further designs a ProbSparse self-attention only using top queries with the biggest measurement results, which can reduce the quadratic complexity in time and memory.
-Pyraformer [153] constructs a multi-resolution -ary tree and develops a Pyramidal Attention Mechanism, in which every node can only attend to its neighboring, adjacent, and children nodes.
-With the calculated mask for attention, Pyraformer can capture both short- and long-temporal dependencies with linear time and space complexity.
+Informer [21] introduces a Query Sparsity Measurement, where a larger value indicates a higher chance of containing the dominant information in self-attention.
+Building on this, it designs ProbSparse self-attention, which selectively uses only the top queries with the highest measurement results, thereby significantly reducing quadratic complexity in both time and memory.
+Pyraformer [154] constructs a multi-resolution tree and develops a Pyramidal Attention Mechanism, in which every node can only attend to its neighboring, adjacent, and children nodes.
+With a calculated mask for attention, Pyraformer captures both short- and long-temporal dependencies with linear computational complexity.
 4.5.2 Patch-wise Dependency
 Patch-based architectures play a crucial role in the Transformer models for both Natural Language Processing (NLP) [16] and Computer Vision (CV) [19].
-Since point-wise representations are insufficient to capture local semantic information in temporal data, several studies [23, 154, 155] have been devoted to exploring patch-level temporal dependencies within time series data.
-Pioneer work Autoformer [22] proposes an Auto-Correlation Mechanism, which captures the series-wise dependencies of time series to replace canonical point-wise self-attention.
+Since point-wise representations often fall short in capturing local semantic information in temporal data, several studies [23, 156, 157] have been devoted to exploring patch-level temporal dependencies in time series data.
+Pioneer work Autoformer [22] proposes an Auto-Correlation Mechanism to replace canonical point-wise self-attention, enabling the model to learn series-wise dependencies.
 Based on the stochastic process theory [158], Auto-Correlation utilizes the Fast Fourier Transform to discover the time-delay similarities between different sub-series.
-A time delay module is further proposed to aggregate the similar sub-series from underlying periods instead of the relation between scattered points, which firstly explores the sub-series level modeling in Transformer-based models.
-Different from modifying the attention mechanism, most recent works utilize patch-wise representations of time series data and perform a self-attention mechanism to capture patch-wise dependencies [23, 54, 157].
-PatchTST [23] and follow-up works split time series into a sequence of overlapped patches and embed each patch following:
-Assume , is patch length and the corresponding number of patches split, and denotes the -th patch with sequence length .
+A time delay module is further proposed to aggregate similar sub-series from underlying periods instead of the relation between scattered points, which firstly explores the sub-series level modeling in Transformer-based models.
+Different from modifying the attention mechanism, most recent works utilize patch-wise representations of the data and perform a canonical self-attention mechanism to capture patch-wise dependencies [23, 52, 159].
+PatchTST [23] and its subsequent works split time series into a sequence of overlapped patches and embed each patch following:
+Let denote the patch length and the corresponding number of patches, and is the -th patch of length .
 The patches are mapped to the latent space through a temporal linear projection and a learnable position embedding .
 Based on the vanilla attention mechanism, PatchTST [23] learns the patch-wise dependencies.
-Going beyond PatchTST, recent Pathformer [158] proposes a multi-scale Transformer-based model with adaptive pathways.
-Based on the patch division of different scales, the adaptive pathways select the patch sizes with the top weights generated by the router to capture multi-scale characteristics.
-The success of PatchTST also benefits from channel-independence design, where each temporal patch-level token only contains information from a single series.
+Going beyond PatchTST, Pathformer [160] introduces a multi-scale Transformer that employs adaptive pathways to dynamically select patch sizes, thereby learning multi-scale temporal representations.
+The success of PatchTST also benefits from its channel-independence design, where each temporal patch-level token encapsulates information from only a single series.
 In addition to capturing the patch-level temporal dependencies within one single series, recent approaches [52, 159] have endeavored to capture interdependencies among patches from different variables over time.
@@
 4.5.3 Series-wise Dependency
-Further expanding the receptive field, there are also some works that attempt to use the tokenization of the whole time series to capture inter-series dependencies.
-iTransformer [30] introduce to multivariate data, and for -th variable , it can be simply formulated as follows:
+Further expanding the receptive field, recent works tokenize the entire time series to capture inter-series dependencies.
+iTransformer [39] introduces for multivariate data and for -th variable , it can be formulated as:
 where is instantiated as trainable linear projector.
@@
 5 Time Series Library
-Time series analysis has emerged as an important research area, attracting significant attention from both academia and industry.
-Recently, extensive exploration of deep learning based methods for time series analysis has resulted in significant advances.
-However, the issue of fair benchmarking poses a pressing challenge in this domain.
-The absence of fair, rational, and comprehensive benchmarks can lead to biased comparisons between different methods and hinder accurate evaluation of their effectiveness, potentially inflating domain advances or hindering practical applications.
-This presents a substantial obstacle to understanding advances and fostering robust development within the field.
-In the domain of time series analysis, several benchmarks have been proposed, such as DGCRN [160], LibCity [161], DL-Traff [162], TS-bench [163], and BasicTS [14].
-More specifically, Autoformer [22] proposed a standard long-term forecasting benchmark covering different practical applications.
+Recently, extensive exploration of deep time series methods has resulted in significant advances.
+Given such booming development, the issue of fair benchmarking poses a pressing challenge for future research.
+Thus, focusing on time series analysis, several benchmarks have been proposed, such as DGCRN [162], LibCity [163], DL-Traff [164], TS-bench [165], and BasicTS [14].
+Besides, Autoformer [22] proposed a standard long-term forecasting benchmark covering different practical applications.
 Further, to verify the generality of different time series analysis models, TimesNet [3] builds a more comprehensive model generalization benchmark covering five mainstream time series analysis tasks.
@@
 Moreover, these benchmarks often fail to provide detailed discussions and comprehensive summaries of task types, model architectures, and specific baseline methods.
-As a result, they do not effectively guide the design of more efficient time series analysis methods or drive further development in the field.
-To effectively address these issues, we introduce and implement Time Series Library (TSLib), a benchmark for fair and comprehensive comparing and evaluating the performance of deep time series models across various time series analysis tasks.
-As shown in Figure 9, TSLib encompasses a unified model experiment pipeline, standardized evaluation protocols, extensive and diverse real-world datasets, mainstream and advanced time series analysis models, and unified experimental validation and analysis process.
-In our Time Series Library, we meticulously followed the official codes and implemented 24 widely used and advanced deep time series analysis models.
-These models are derived from four canonical deep learning architectures.
+As a result, they do not effectively guide the design of more efficient time series methods or drive further development in the field.
+To address these challenges effectively, we introduce Time Series Library (TSLib), a benchmark designed for the fair and comprehensive evaluation of deep time series models across a wide range of analysis tasks.
+As shown in Figure 8, TSLib provides a unified experimental framework that includes standardized evaluation protocols, a rich and diverse collection of real-world datasets, state-of-the-art and widely-used time series models, and a streamlined process for experimental validation and analysis.
+Although numerous time series libraries have been proposed in recent years, TSLib stands out by including advanced models and supporting a wide range of time-series analysis tasks, which have earned widespread recognition from the community as shown in Table III.
+Specifically, we meticulously followed the official codes and implemented 30 widely used and advanced deep time series models.
 Users can choose from these models based on their specific practical usage scenarios.
 The code is available at https://github.com/thuml/Time-Series-Library.
-As follows, we will provide a detailed description of our TSLib, including the design and implementation principles (Section 5.1), the evaluation protocols and metrics (Section 5.2), the dataset descriptions (Section 5.3), and the main results of models with different architectures (Section 5.4).
+Library |
+Models |
+Task |
+Stars |
+tslearn∗ [166] |
+10 |
+Clustering, Classification, Regression |
+3k |
+StatsForecast∗ [167] |
+31 |
+Forecasting, Anomaly Detection |
+4.4k |
+NeuralForecast [168] |
+33 |
+Forecasting |
+3.6k |
+GluonTS [169] |
+25 |
+Forecasting, Anomaly Detection |
+4.9k |
+| TSLib (This Survey) | 30 |
+Forecasting, Imputation |
+9.4k |
+Classification, Anomaly Detection |
+This section provides an overview of TSLib.
+We will introduce design and implementation principles of TSLib in Section 5.1, followed by a brief description of default evaluation metrics in Section 5.2, supported datasets in Section 5.3, and experimental settings in Section 5.4.
+Finally, comparisons of various models are included in Section 5.5.
 5.1 Design and Implementation Principle
-TSLib is designed based on the well-established factory pattern and implements a unified interface between data and model objects, thus enabling a clear separation between deep model creation and usage, promoting modularity and flexibility.
-By loading different data and model objects and combining specific task heads during model training, TSLib enables different datasets and models to be shared and extended, allowing easy switching between various time series analysis tasks.
-These design and implementation principles provide enhanced flexibility and scalability for our TSLib.
-Furthermore, as illustrated in Figure 9, TSLib introduces a unified experimental pipeline covering the overall process of the model training and evaluation, which includes data source, data processing, model training and analysis, and model performance evaluation.
-Data Source Our TSLib provides extensive support for a wide range of diverse and multi-type datasets in a variety of formats, including ”.csv”, ”.npz”, ”.txt”, etc.
-As shown in Figure 9 and Table III, TSLib currently supports more than 30 datasets with different sampled frequencies across four mainstream time series analysis tasks, all derived from real-world scenarios in domains such as energy, transportation, economics, weather, and medicine, etc.
-Moreover, TSLib excels in scalability, allowing for the effortless integration of new data sources of different data types.
-Data Processing Data processing plays a pivotal role in guaranteeing stable training within the realm of time series analysis.
-Within the Time Series Library, a multitude of data processing steps are conducted, including time window splitting, data batch generation, etc.
-Subsequently, the raw data is partitioned into separate sets for training, validation, and testing purposes, enabling streamlined model training and equitable comparisons.
+TSLib is designed based on the well-established factory pattern and implements a unified interface between data, model, and experiment objects, thus enabling a clear separation between deep model creation and usage, promoting modularity and flexibility.
+By loading different data and model objects and combining task-specific heads, TSLib enables different datasets and models to be shared and extended, allowing easy switching between various datasets, models, and analysis tasks.
+Here are the key features.
+Diverse Data Formats TSLib natively supports a wide range of datasets in a variety of formats, including “.csv”, “.npz”, “.txt”, etc.
+As summarized in Figure 8, TSLib currently supports more than 30 datasets with different sampled frequencies across four mainstream time series analysis tasks, all derived from real-world scenarios, such as energy, transportation, economics, weather, and medicine etc.
+Moreover, TSLib excels in scalability, allowing for the effortless integration of new data sources of different types.
+Besides, to support subsequent training and evaluation, TSLib includes a multitude of data processing steps, including time window splitting, data batch generation, etc, where the raw data is partitioned into separate windows for training, validation, and testing purposes, enabling streamlined model training and equitable comparisons.
 These steps serve as indispensable prerequisites for attaining precise and dependable results across a range of diverse time series analysis tasks.
-Moreover, our TSLib provides additional support for numerous crucial and effective data processing strategies based on different model design principles [39], [40], [41], [45] to enhance model performance and training efficiency.
-We encapsulate these various design strategies within our basic data processing layer, encompassing techniques such as data normalization, time-frequency decomposition, Fourier analysis, and more.
-When utilizing TSLib, users have the flexibility to select these strategies to improve training effect based on their specific requirements and objectives.
-Model Training and Analysis After the data processing phase, the raw time series data is transformed into the desired format for model training.
-Model training forms the crux of the entire experiment pipeline, where we fine-tune the model parameters based on the input to predict the output with minimal error.
-Our primary goal during model training is to obtain the best possible trainable parameters that result in a significant improvement in model performance.
-Each model has its own unique design and training objective.
-The model analysis procedure is to determine the optimal model parameters by comparing the correlation between the training and validation losses.
-Our TSLib includes complete log printing and result storage functions record and evaluate the training process.
-By employing rational model analysis techniques, we can efficiently obtain models with superior performance and stronger generalization.
-Performance Evaluation Model evaluation is a crucial step in verifying the effectiveness and generalization of trained time series models.
-It involves model prediction and performance evaluation, providing insights into the efficacy of the trained model.
-In TSLib, we provide evaluation support for four mainstream time series analysis tasks: classification, imputation, forecasting (long-term or short-term), and anomaly detection.
-Each task comes with its specific evaluation metric, enabling a comprehensive assessment of the performance of models.
-These metrics play a crucial role in determining the effectiveness of the trained model and its suitability for the intended task.
-| Task | Dataset | Dimension | Length | Domain | Size |
-| Classification | EthanolConcentration | 3 | 1,751 | Alcohol Industry | 20.3 MB |
-| FaceDetection | 144 | 62 | Face (250 Hz) | 789.1 MB | |
-| Handwriting | 3 | 152 | Motion | 3.9 MB | |
-| Heartbeat | 61 | 405 | Health (0.061 secs) | 87.8 MB | |
-| JapaneseVowels | 12 | 29 | Voice | 1.1 MB | |
-| PEMS-SF | 963 | 144 | Transportation (1 day) | 420.1 MB | |
-| SelfRegulationSCP1 | 6 | 896 | Health (256 Hz) | 17.8 MB | |
-| SelfRegulationSCP2 | 7 | 1,152 | Health (256 Hz) | 17.7 MB | |
-| SpokenArabicDigits | 13 | 93 | Voice (11025 Hz) | 37.6 MB | |
-| UWaveGestureLibrary | 3 | 315 | Gesture | 3.4 MB | |
-| Imputation | ETTh1, ETTh2 | 7 | 17,420 | Electricity (1 hour) | 10.4 MB |
-| ETTm1, ETTm2 | 7 | 69,680 | Electricity (15 mins) | 2.6 MB | |
-| Electricity | 321 | 26,304 | Electricity (1 hour) | 95.6 MB | |
-| Weather | 21 | 52,696 | Environment (10 mins) | 7.2 MB | |
-| Long-term Forecasting | ETTh1, ETTh2 | 7 | 17,420 | Electricity (1 hour) | 10.4 MB |
-| ETTm1, ETTm2 | 7 | 69,680 | Electricity (15 mins) | 2.6 MB | |
-| Electricity | 321 | 26,304 | Electricity (1 hour) | 95.6 MB | |
-| Weather | 21 | 52,696 | Environment (10 mins) | 7.2 MB | |
-| Traffic | 862 | 17,544 | Transportation (1 hour) | 136.5 MB | |
-| Exchange | 8 | 7,588 | Economic (1 day) | 623 KB | |
-| ILI | 7 | 966 | Health (1 week) | 66 KB | |
-| Short-term Forecasting | M4-Yearly | 1 | 6 | Demographic | 589.5 MB |
-| M4-Quarterly | 1 | 8 | Finance | ||
-| M4-Monthly | 1 | 18 | Industry | ||
-| M4-Weakly | 1 | 13 | Macro | ||
-| M4-Daily | 1 | 14 | Micro | ||
-| M4-Hourly | 1 | 48 | Other | ||
-| Anomaly Detection | SMD | 38 | 100 | Industry (1 min) | 436.4 MB |
-| MSL | 55 | 100 | Industry (1 min) | 58.2 MB | |
-| SMAP | 25 | 100 | Industry (1 min) | 113.0 MB | |
-| SwaT | 51 | 100 | Industry (1 min) | 903.2 MB | |
-| PSM | 25 | 100 | Industry (1 min) | 107.1 MB |
-5.2 Evaluation Protocols
-In order to conduct a fair and comprehensive model performance verification, our Time Series Library is designed to provide standardized evaluation protocols for four mainstream time series analysis tasks following [3].
-The primary goal of these standardized and unified evaluation protocols is to quantify the effectiveness of different time series analysis methods with varying architectures.
-Additionally, they provide valuable insights into the strengths and limitations of different methods across diverse time series analysis tasks.
-By establishing these standardized evaluation protocols, we aim to promote fair comparisons between different methods and improve our understanding of their performance in various time series analysis scenarios.
-For long-term forecasting and imputations, we rely on Mean Square Error (MSE) and Mean Absolute Error (MAE) as the primary evaluation metrics.
-These metrics help us accurately assess the accuracy of our predictions and imputations.
+Moreover, TSLib provides additional support for numerous crucial and effective data processing strategies [36, 37, 38, 43] to enhance model performance and training efficiency.
+We encapsulate these various design strategies as basic data processing layers, encompassing techniques such as data normalization, time-frequency decomposition, Fourier analysis, and more.
+When utilizing TSLib, users have the flexibility to select these strategies to improve the training effect based on their specific requirements and objectives.
+Flexible Model Construction As stated in the previous sections, we abstract the construction of deep time series models as the combination of basic modules (Section 2) and model architectures (Section 4).
+Specifically, we separate the most commonly used modules, such as stationarization and decomposition, from previous elaborately designed models.
+In this way, it is very convenient to reproduce or add a deep model with pre-defined basic modules.
+Further, although different models are originally proposed for different tasks, such as Autoformer [22] is for forecasting and TimesNet [3] is for more generic tasks, following the “model-centric” principle in this survey, we unify diverse analysis tasks as a representation learning task with domain-specific heads, enabling a task-general evaluation for different backbones.
+Comprehensive Performance Evaluation Building upon the above two essential features, TSLib can naturally support the evaluation of diverse models on different tasks based on extensive datasets, building a comprehensive benchmark.
+Specifically, in TSLib, we provide evaluation for four mainstream time series analysis tasks: classification, imputation, forecasting, and anomaly detection.
+Each task involves specific evaluation metrics, enabling a comprehensive assessment of the performance of models.
+Noticing that, we further distinguish the forecasting task as long- and short-term tasks following the previous convention, where the separation depends on the forecasting horizon.
+5.2 Evaluation Protocols and Metrics
+To conduct a fair and comprehensive model comparison, TSLib provides standardized evaluation protocols for four mainstream time series analysis tasks following [3].
+Specifically, different models are trained and tested based on the same training strategies and test sets to avoid potential unaligned configurations and ensure a rigorous comparison among different models.
+Following previous research, for long-term forecasting and imputations, we adopt Mean Square Error (MSE) and Mean Absolute Error (MAE) as the primary evaluation metrics.
 For short-term forecasting, we use the Symmetric Mean Absolute Percentage Error (SMAPE) and Mean Absolute Scaled Error (MASE) as metrics, which focus on absolute errors and reduce the impact of outliers, providing reliable evaluations of forecast accuracy across different datasets and methodologies.
-In the case of time series classification tasks, we utilize Accuracy as the evaluation metric.
-Accuracy measures the overall prediction performance by calculating the ratio of correctly classified samples to the total number of samples.
+In time series classification, we utilize classification accuracy as the evaluation metric.
 For anomaly detection, we employ the F1 score to validate the identification of abnormal values.
-The F1 score represents a balanced combination of precision and recall, offering a comprehensive assessment of a classifier’s performance, especially when dealing with imbalanced classes in the context of anomaly detection.
-5.3 Datasets
-TSLib includes a variety of mainstream datasets across different numbers of samples and categories, a richness of tasks, and a diversity of domains.
-In this section, we will focus on introducing representative datasets from various time series analysis tasks included in the Time Series Library.
-Classification Time series classification aims to assign a label or category to a time series based on its temporal features.
-To evaluate this capability, we selected ten multivariate datasets from the UEA Time Series Classification Archive [164], supported in our TSLib.
-These datasets cover a range of practical tasks, including gesture, action, audio recognition, and medical diagnosis through heartbeat monitoring.
-We pre-processed the datasets according to the descriptions provided in [165].
-Detailed dataset descriptions are shown in Table III.
+The F1 score represents a balanced combination of precision and recall, offering a comprehensive assessment of a classifier’s performance, especially when dealing with imbalanced abnormal classes.
+5.3 Supported Datasets
+TSLib natively supports a variety of widely used datasets across different numbers of samples and variables, covering diverse domains and tasks.
+The correspondences between tasks and datasets are depicted in Figure 8.
+Since all datasets are commonly used, here is only a brief introduction, and more details can be found in previous papers [3].
+Forecasting Forecasting is one of the most widely explored tasks in previous research, bringing us a wide choice of experimental datasets.
+Just as stated in design principles, we further divide this task into two types (long- and short-term) according to the forecasting horizon.
+For long-term time series forecasting, TSLib includes nine datasets, including Electricity Transformer Temperature (ETT, including four subsets) [21], Electricity, Weather, Traffic, Exchange, and Illness [22].
+For the short-term forecasting task, we include the M4 dataset [170], which comprises six sub-datasets with varying sampling frequencies and domains.
 Imputation Due to glitches, the collected time series data may contain partially missing values, posing a challenge for time series analysis.
-Therefore, time series imputation is a crucial task in real-world applications, aiming to fill in missing values within a time series based on contextual observations from the data.
-For our benchmark, we selected Electricity Transformer Temperature (ETT) [21], Electricity, and Weather to evaluate the performance of time series imputation tasks with different missing ratios.
-Forecasting Time series forecasting is an essential task in time series analysis, and it has been widely explored in academic and industry domains.
-By leveraging historical patterns and trends, the model can predict future values or trends of a time series.
-Time series forecasting can be broadly divided into two types: long-term forecasting and short-term forecasting.
-For the long-term time series forecasting task, a wide range of datasets are included in our benchmark, including Electricity Transformer Temperature (ETT), Electricity, Weather, Traffic, Exchange, and Illness (ILI).
-For the short-term forecasting task, we selected the M4 dataset [166], which comprises six sub-datasets with varying sampling frequencies and domains.
+However, in real-world applications, it is really hard to obtain the ground truth values for the missing records.
+Thus, we utilize the Electricity Transformer Temperature (ETT, with four subsets) [21], Electricity, and Weather to construct the missing value situation by randomly masking a certain portion of time steps.
 Anomaly Detection Anomaly detection involves identifying unusual or abnormal patterns in a time series.
@@
 There are some mainstream anomaly detection datasets supported in TSLib, such as Server Machine Dataset (SMD) [171], Mars Science Laboratory rover (MSL) [172], Soil Moisture Active Passive satellite (SMAP) [172], Secure Water Treatment (SWaT) [173], and Pooled Server Metrics (PSM) [174] which are collected from a variety of industrial scenarios.
-5.4 Main Results
-To examine the strengths and limitations of various methods in mainstream time series analysis tasks, we select twelve representative models from our TSLib.
-These models encompass four popular deep model architectures and address tasks, including long-term and short-term forecasting, classification, imputation, and anomaly detection.
-Baselines To conduct a fair comparative analysis and thoroughly explore the effectiveness of different model architectures in various time series analysis tasks, we conduct comparative comparison experiments using state-of-the-art models designed based on different deep architectures.
+Classification We selected ten multivariate datasets from the UEA Time Series Classification Archive [175] to test the series-level classification tasks.
+These datasets cover a range of practical tasks, including gesture, action, audio recognition, and medical diagnosis through heartbeat monitoring.
+We pre-processed the datasets according to the descriptions in [176].
+5.4 Default Experimental Setups
+Here, we provide experimental setups for each analysis task supported by TSLib.
+For long-term forecasting, we use a look-back length of 36 and prediction lengths of for the ILI dataset.
+For other datasets, a unified look-back length of 96 is adopted, with forecast horizons of .
+In short-term forecasting, the prediction lengths are set as for yearly, quarterly, monthly, weekly, daily, and hourly data, respectively.
+In the classification task, we aim to categorize high-dimensional time series data into multiple classes.
+In the imputation task, we randomly mask the time points in ratios of .
+In the anomaly detection task, the reconstruction is a classical task for unsupervised point-wise representation learning, where the reconstruction error is a natural anomaly criterion.
+Therefore, we set deep time series models as the base models for reconstruction and utilize the reconstruction error as a shared anomaly criterion across all experiments.
+| Task | LR∗ | Loss | BatchSize | Epochs |
+| Long-term Forecasting | MSE | 32 | 10 | |
+| Short-term Forecasting | SMAPE | 16 | 10 | |
+| Imputation | MSE | 16 | 10 | |
+| Classification | Cross Entropy | 16 | 100 | |
+| Anomaly Detection | MSE | 128 | 10 |
+Implementation Details In this paper, all experiments were implemented using PyTorch and executed on an NVIDIA A100 GPU.
+The Adam optimizer was used for model optimization.
+Each model’s performance was comprehensively evaluated across various time series analysis tasks.
+To achieve optimal performance and ensure fair comparisons across models and datasets, we conducted hyperparameter searches for each model architecture following [177].
+Specifically, we search the number of hidden layers and the dimension of hidden representations from the range .
+For each time series analysis task, TSLib utilized task-specific and model-shared training configurations, including learning rate, loss function, batch size, and number of epochs, which are listed in Table IV.
+Look-back Searching Setting Note that in long-term forecasting tasks, the model’s performance is highly dependent on the historical length of the input series.
+Given the inconsistent look-back length of various papers, we further conducted an additional set of experiments referred to as “Look-back Searching” to ensure a fair and reliable comparison.
+In these experiments, we systematically searched for input length among following [177] and performed an extensive hyperparameter search on each model’s architecture to reach their optimal performance.
+This approach allowed us to assess the maximum achievable performance of each model while ensuring that the reported results accurately reflect their capabilities and facilitate a fair comparison across different models.
+5.5 Model Analysis
+To examine the strengths and limitations of various methods in mainstream time series analysis tasks, we select thirteen representative models from TSLib for comparison.
+These models encompass five popular deep model architectures and tasks, including long-term and short-term forecasting, classification, imputation, and anomaly detection.
+Baselines To thoroughly explore the effectiveness of different model architectures in various tasks, we conduct comparative experiments using state-of-the-art models designed based on different deep architectures.
 As shown in Figure 9, we select several advanced and representative Transformer-based models: iTransformer [39], PatchTST [23], Autoformer [22], Non-Stationary Transformer (Stationary) [38], and FEDformer [83] to verify the performance.
@@
 For RNN-based models, we included the novel and effective Mamba [118].
-Finally, we include DLinear [52], N-BEATS [24], and TiDE [97] as representative MLP-based models for analysis.
-It is important to mention that TiDE [97] is designed to be dependent on specific timestamps, and cannot be easily adapted to part tasks without timestamps, such as short-term forecasting, anomaly detection, and classification.
-Unified Experimental Settings For the long-term forecasting task, we conducted two experimental settings to ensure a fair and comprehensive comparison: unified hyperparameter and hyperparameter searching.
-For the unified hyperparameter, we conducted experiments using a range of standardized hyperparameters set across different datasets.
-This allows us to accurately evaluate the relative performance of time series models with different deep architectures while keeping other factors constant.
-As for the “hyperparameter searching” scenario, we conducted separate hyperparameter searches for different model architectures and time series datasets.
-This approach enabled us to identify the best performance of different time series analysis models.
-By employing above both settings, we obtain a comprehensive understanding of the forecasting performance of different time series models.
-For the remaining tasks, we maintained the standard experimental settings as outlined in [3] to validate the performance of different time series models.
-Overall Results Based on the overall results of the models across different architectures in Figure 11, we surprisingly find that the MLP-based models, which are generally simpler and have lower computational overhead, perform well on the time series forecasting task.
-However, these models appear to be less effective in other types of tasks, which requires the model to learn more informative representations.
+As for GNN-based models, since most approaches require a dependency graph as priors for the model, we take StemGNN as the representative model, which enables a data-driven construction of dependency graphs.
+DLinear [50], N-BEATS [24], and TiDE [96] are tested as representative MLP-based models.
+Since TiDE [96] depends on timestamps and cannot be easily adapted to tasks without timestamps in their datasets, such as short-term forecasting, anomaly detection, and classification, we only test it on the long-term forecasting task.
+Overall Results As shown in Figure 9, we can find that the Transformer-based models, e.g., iTransformer [39] and PatchTST [23], exhibit superior forecasting capabilities compared to other models for both long-term and short-term forecasting tasks.
+This further proves that it is of great significance to explore time series tokenization strategies, which is the key to unlocking the capacity of Transformer-based models.
+Additionally, TimesNet [3] shows a more comprehensive and effective performance covering time series classification, imputation, and anomaly detection tasks, pioneering a milestone in task-general time series models.
+Beyond detailed comparison for each specific model, we also include a summarized performance comparison among different types of deep learning backbones in the left of Figure 10, where the performance of each type of backbone is the best performance achieved among all corresponding models.
+We surprisingly find that the MLP-based models, which are generally simpler and have lower computational overhead, perform well on time series forecasting.
+However, these models appear to be less effective in other types of tasks, which require the model to learn more informative representations.
 On the contrary, the CNN-based models demonstrate more comprehensive capabilities and excel in classification, imputation, and anomaly detection tasks.
@@
 It further shows that Transformer-based models hold significant research value and application potential in the field of time series analysis and have emerged as a particularly promising option in the time series domain.
-As illustrated in Figure 10, we have also included more detailed results and a top three performance leaderboard for four representative time series analysis tasks.
-These results clearly show that the Transformer-based models, namely iTransformer [30] and PatchTST [23], exhibit superior forecasting capabilities compared to other models for both long-term and short-term forecasting tasks.
-This further proves that it is of great significance and value to explore different modeling methods of temporal tokens in time series.
-Additionally, TimesNet [3] shows a more comprehensive and effective performance covering time series classification, imputation, and anomaly detection tasks.
-It has pioneered a milestone in the general time series analysis model.
-We believe that TSLib can provide useful start code, valuable insights on model properties and model selection guidance for future research and real-world applications.
+We believe that TSLib can provide useful start code, valuable insights on model properties, and model selection guidance for future research and real-world applications.
+Efficiency Analysis We also provide an efficiency comparison to offer deeper insights into the practicality and feasibility of different approaches.
+Specifically, we measured the training time, inference time, and memory usage of each model in performing long-term forecasting, where all models operated under unified hyperparameter settings.
+As shown in Figure 10, it is observed that the tokenization strategy also affects the model efficiency, where Autoformer [22] with point-wise tokens is slower than PatchTST [23] with patch-wise tokens and iTransformer [39] with series-wise tokens.
+Besides, CNN-based models, such as SCINet [130] and TimesNet [3], are much more time-consuming compared to other architectures since they have to adopt diverse kernels to capture intricate temporal patterns.
+In contrast, MLP-based models and the latest Mamba model stand out for their superior efficiency.
+Their simpler and more lightweight architectures make them particularly suitable for scenarios with limited computational resources.
+The above findings underscore the inherent trade-offs between model efficiency and predictive performance, which could also be a valuable guide for selecting models for resource-constrained scenarios.
 6 Future Directions
-In this section, we provide a discussion on the promising directions for time series analysis.
+In this section, we present a discussion on promising directions for time series analysis.
 6.1 Time Series Pre-training
-A pretraining-finetuning learning paradigm is a two-stage approach commonly used in nature language processing (NLP) [171], [16], [172], [173], [173] and computer vision (CV) [174], [175], [176], [177].
+The pretraining-finetuning learning paradigm is a two-stage approach commonly used in nature language processing (NLP) [178], [16], [179], [180], [180] and computer vision (CV) [181], [182], [183], [184].
 Pre-training establishes the basis of the abilities of Large Models through unsupervised learning [185, 186].
 Fine-tuning can improve the performance of the pre-trained model on a specific task or domain.
-Due to the limited availability of labeled datasets, self-supervised pre-training [175] has garnered significant attention and has been extensively investigated in the domains of natural language modeling and computer vision.
-Self-supervised pre-training paradigm significantly reduces labeling expenses and benefits for diverse downstream tasks.
-Notably, recent research efforts have introduced several self-supervised pre-training methods tailored for time series data, which can be primarily classified into contrastive learning [180] and masked time series modeling [16, 181].
-Contrastive learning refers to learning the representations of data by contrasting between similar and dissimilar pairs, where similar sample pairs are learned to be close to each other and dissimilar pairs are far apart [180, 182, 183].
+Given the scarcity of labeled datasets, self-supervised pre-training [182] has garnered significant attention and has been extensively investigated in the domains of natural language modeling and computer vision.
+This paradigm significantly reduces labeling expenses and yields considerable benefits for diverse downstream tasks.
+Notably, recent research has extended self-supervised pre-training methods to time series data, primarily categorizing them into contrastive learning [187] and masked time series modeling [16, 188].
+Contrastive learning refers to learning the representations of data by contrasting similar and dissimilar pairs, where similar sample pairs are learned to be close to each other and dissimilar pairs are far apart [187, 189, 190].
 Although SimCLR [189] has demonstrated remarkable success in the domain of computer vision, directly applying SimCLR to the field of time series data yields unsatisfactory results due to the insufficient modeling of temporal dependencies.
-CPC [184] introduced contrastive predictive coding, which utilizes model-predicted features as positive samples and randomly-sampled features as negative samples, to obtain time series representations that are advantageous for downstream tasks.
-TimeCLR [185] proposes a DTW data augmentation to generate phase shift and amplitude-change phenomena which can preserve time series structure and feature information.
+CPC [191] introduces contrastive predictive coding, which utilizes model-predicted features as positive samples and randomly-sampled features as negative samples, to obtain time series representations that are advantageous for downstream tasks.
+TimeCLR [192] proposes a DTW data augmentation to generate phase shift and amplitude-change phenomena, which can preserve time series structure and feature information.
 TS-TCC [193] employs efficient data augmentations designed for time-series data, and learns discriminative representations from the proposed Temporal Contrasting module and Contextual Contrasting module.
@@
 CoST [197] proposes a contrastive learning framework to learn disentangled seasonal-trend representations for long sequence time series data.
-TF-C [191] develop frequency-based contrastive augmentation to leverage rich spectral information and explore time-frequency consistency.
+TF-C [198] develops frequency-based contrastive augmentation to leverage rich spectral information and explore time-frequency consistency.
 Masked modeling is a reconstruction-based method, which can predict a masked token in a sequence based on the context unmasked part [199, 16, 188].
@@
 Further, PatchTST [23] segments the time series data into multiple non-overlapping patches and proposes a patch-level masked modeling approach.
-SimMTM [194] relates masked modeling to manifold learning and presents a neighborhood aggregation design for reconstruction based on the similarities learned in series-wise representation space.
+SimMTM [201] relates masked modeling to manifold learning and presents a neighborhood aggregation design for reconstruction based on the similarities learned in the series-wise representation space.
 HiMTM [202] proposes a novel hierarchical masked time series pre-training framework to effectively capture the multi-scale characteristics of time series data.
@@
 Nevertheless, in real-world forecasting scenarios, it is crucial to involve additional information that is related to the temporal variation of the target time series and must be taken into account, such as weather conditions or holidays.
-MOIRAI [201] tries to flatten multivariate time series into a single sequence containing all variate, but its generalization capabilities to other downstream analysis tasks are under-explored.
-In addition to modeling inter-series dependencies, modeling the relationship between time series and external factors in other can achieve a better understanding of time series data.
-These external factors may be in the form of other modalities, such as text data or calendar data, and thus multimodal learning is a future trend in the development of a multi-modal time series foundation model.
+MOIRAI [208] tries to flatten multivariate time series into a single sequence containing all variates, but its generalization capabilities to other downstream analysis tasks are under-explored.
+Beyond merely modeling inter-series dependencies, incorporating the relationship between time series and external factors offers a deeper understanding of the data.
+These external factors can manifest as other modalities, such as textual information or calendar data.
+Consequently, multimodal learning represents a promising direction for future time series foundation models.
 6.2.2 Adaptation of Large Language Models
@@
 Benefiting from the generative capability of LLMs, most of the existing research focuses on time series forecasting tasks.
-GPT4TS [207] propose a unified framework for diverse time series analysis tasks by using a pre-trained GPT-2 model and fine-tuning the positional embeddings and the parameters of the layer normalization for each analysis task.
+GPT4TS [156] proposes a unified framework for diverse time series analysis tasks by using a pre-trained GPT-2 model and fine-tuning the positional embeddings and the parameters of the layer normalization for each analysis task.
 Prompting Large Language Models Recent Large Language Models exhibit the strong abilities of in-context learning [147] and instruction following [214].
@@
 There are other works represented by Autotimes [53, 221], that attempt to design soft prompts for time series data.
-However, existing prompting approaches are tailored for forecasting, and how to empower LLMs to other time series tasks besides forecasting is relatively unexplored.
+However, existing prompting approaches are tailored for forecasting, and how to empower LLMs to perform other time series tasks besides forecasting is relatively unexplored.
+Both of the above two directions would require a large-scale training dataset.
+However, due to the privacy of industrial or IoT scenarios, the accessible time series data is limited, necessitating the exploration of time series generation techniques [222, 213].
+Generating high-quality and diverse time series data not only provides valuable resources for model training but also holds the potential to significantly enhance the performance, robustness, and generalization capabilities of these models [223].
 6.3 Practical Applications
+6.3.1 Probabilistic Forecasting
+Probabilistic forecasting is a pivotal advancement in practical applications of deep time series models, particularly for informed decision-making.
+Unlike previous widely-studied deterministic forecasting, probabilistic forecasting offers a distribution over potential future outcomes, enabling stakeholders to assess risks, evaluate confidence intervals, and make robust decisions under uncertainty.
+Most existing probabilistic forecasting models rely on predefined probability distributions and are designed to predict the parameters of the distributions in the future.
+While these methods have achieved significant success, the dependence on predefined distributions can limit their ability to represent complex and heterogeneous real-world uncertainties.
+Therefore, accurately modeling and representing such complexities to generate more flexible and robust probabilistic forecasts remains a critical and active area of future research.
 6.3.2 Handling Extremely Long Series
-Deep time series models have demonstrated remarkable performance across a wide range of downstream tasks, their applicability to longer time series data is often limited by scalability and high computational complexity.
+Deep time series models have demonstrated remarkable performance across a wide range of downstream tasks, but their applicability to longer time series data is often limited by scalability and high computational complexity.
 In industrial time series analysis, high-frequency sampling results in lengthy historical data, impeding the practical implementation of advanced deep models.
@@
 However, in practical applications, different from multivariate time series analysis, the main variables and covariates usually occupy different positions.
-Given the crucial role played by exogenous variables in real-world applications, it is essential to explore a unified framework for modeling the relationships between the endogenous and exogenous variants, which allows for a more comprehensive understanding of interrelations and causality among different variants, leading to better and more reliable model performance, as well as interpretability.
+Given the crucial role played by exogenous variables in real-world applications, it is essential to explore a unified framework for modeling the relationships between the endogenous and exogenous variables, which allows for a more comprehensive understanding of interrelations and causality among different variants, leading to better and more reliable model performance, as well as interpretability.
 6.3.4 Processing Heterogeneous Data
@@
 This may involve developing new techniques for feature extraction, incorporating domain knowledge into model design, or exploring alternative architectures that are better suited to handle variable-length inputs.
-As researchers continue to explore this unexplored area in time series analysis, there is potential for significant advances in areas such as finance, healthcare, and environmental monitoring.
-By improving our ability to model and analyze heterogeneous time series data, we can gain deeper insights into complex systems and make more informed decisions based on predictive analytics.
-Overall, further research in this area holds great promise for advancing our understanding of temporal data dynamics and enhancing the capabilities of time series modeling in real-world applications.
+As researchers continue to explore this relatively uncharted area in time series analysis, there is potential for significant advances in areas such as finance, healthcare, and environmental monitoring.
+Enhancing models’ ability to analyze heterogeneous time series data can lead to deeper insights into complex systems, enabling more informed decision-making through predictive analytics.
+Overall, further research in this area holds great promise for advancing our understanding of temporal data dynamics and significantly enhancing time series modeling capabilities in real-world applications.
 7 Conclusion
-In this survey, we provide a systematic review of deep models in time series analysis and introduce Time-Series Library (TSLib) as a fair benchmark for deep time series models across various analysis tasks.
+In this survey, we provide a systematic review of deep models in time series analysis and introduce Time-Series Library (TSLib) as a fair benchmark for deep time series models across various tasks.
 Compared with previous reviews that focus on a specific analysis task or model architecture, this paper provides a comprehensive survey and overview of existing deep models for time series analysis, ranging from forecasting, classification, imputation, and anomaly detection.
-We first present a detailed review of the universal modules that are widely used among time series models, including normalization, decomposition, and Fourier analysis.
-Next, we summarize existing deep time series models from the perspective of backbone architecture.
-Based on the review of existing literature, we introduce a practical open-source library, Time Series Library (TSLib), which has included representative deep time series models that can be a fair evaluation benchmark in the field of time series analysis.
-Finally, we discuss future research directions for deep time series models based on the recent development of the AI community and the practical application needs of time series analysis in real-world scenarios.
+By separating basic modules from complicated and diverse deep time series models, we successfully uncover the fundamental design principles for current research, forming an inspiring model-centric perspective.
+Besides, based on TSLib, which enables flexible and comprehensive model evaluation, we conduct extensive experiments to bring insights for model design and practical usage.
+With clear summarization, rich experiments, and in-depth discussion, we believe that this survey and TSLib can be instructive and useful for future research on deep time series models.